### v2

# Fuzzy dark matter simulations

Abstract
Fuzzy dark matter (FDM), composed of ultralight bosons, exhibits intricate wave phenomena on galactic scales. Compared to cold dark matter, FDM simulations are significantly more computationally demanding due to the need to resolve the de Broglie wavelength and its rapid oscillations. In this review, we first outline the governing equations and distinctive features of FDM. We then present a range of numerical algorithms for both wave- and fluid-based simulations, discuss their respective advantages and limitations, and highlight representative test problems. To facilitate code comparison, we also provide publicly available initial condition files for both isolated-halo and cosmological simulations.
keywords:
Numerical methods, Dark matter, Large-scale structure, Galaxies
1 Introduction
Fuzzy dark matter (FDM) offers a compelling alternative to cold dark matter (CDM). It consists of ultralight bosons characterized by a single parameter—the boson mass $m_{22}\equiv m/10^{-22}\,\rm eV\sim 1\text{--}10^{3}$. The associated large de Broglie wavelength gives rise to distinctive wave-like phenomena on galactic scales, including the suppression of low-mass halos due to quantum pressure, density fluctuations and vortices from wave interference, and the formation of compact soliton cores as the ground state of the host halo potential. For comprehensive reviews on the theoretical background and observational constraints of FDM, see, for example, Marsh (2016); Niemeyer (2020); Hui (2021); Chavanis (2025); Eberhardt and Ferreira (2025).
This review concentrates on computational aspects of FDM research (for an earlier review, see Zhang et al. 2019). Numerical simulations play an indispensable role in identifying unique FDM properties and confronting model predictions with data. For example, FDM simulations have been used to study the nonlinear evolution of large-scale structure, the halo mass function, the dynamical heating of stellar systems caused by wave interference, and the formation and dynamics of soliton cores. Here we focus on the simplest scenario: a single FDM species without self-interaction, although the simulation techniques discussed can be readily generalized to explore such extensions.
The remainder of this review is organized as follows. First, we introduce the governing equations and general features of FDM in Section 1. We then present a variety of FDM algorithms in Section 2 and describe the associated numerical challenges in Section 3. Representative numerical tests and results are provided in Section 4. Finally, we conclude in Section 5. Links to the initial condition files for the isolated-halo and cosmological simulations are made available in Data availability.
1.1 Governing equations
We present the wave formulation of FDM in Section 1.1.1 and its fluid representation in Section 1.1.2.
1.1.1 Wave formulation
The governing equations of FDM are the Schrödinger–Poisson equations:
$i\frac{\partial\psi}{\partial t}$ $=\left(-\frac{\hbar}{2m}\nabla^{2}+\frac{m}{\hbar}V\right)\psi,$ (1)
$\nabla^{2}V$ $=4\pi G\|\psi\|^{2},$ (2)
where $\psi$ is the wave function, $\hbar$ is the reduced Planck constant, $V$ is the gravitational potential, and $G$ is the gravitational constant. This system of equations is invariant under the following scale transformation (Ruffini and Bonazzola 1969; Guzmán and Ureña-López 2004):
$(m,\bm{r},t,\psi,V)\to(\beta m,\alpha^{-1}\beta^{-1/2}\bm{r},\alpha^{-2}t,\alpha^{2}\psi,\alpha^{2}\beta^{-1}V),$ (3)
where $\alpha$ and $\beta$ are arbitrary dimensionless constants. Note also that $m$ and $\hbar$ always appear as the ratio $m/\hbar$.
For cosmological simulations, one can define $\bm{\tilde{r}}=a^{-1}\bm{r}$, $\tilde{\psi}=a^{3/2}e^{-imHr^{2}/2\hbar}\psi$, and $\tilde{V}=a^{2}(V-V_{b})$ to rewrite Eqs. (12) in comoving coordinates:
$ia^{2}\frac{\partial\tilde{\psi}}{\partial t}$ $=\left(-\frac{\hbar}{2m}\tilde{\nabla}^{2}+\frac{m}{\hbar}\tilde{V}\right)\tilde{\psi}$, (4)
$\tilde{\nabla}^{2}\tilde{V}$ $=4\pi Ga\left(\|\tilde{\psi}\|^{2}-\tilde{\rho}_{\rm b}\right),$ (5)
where $a$ is the scale factor, $H$ is the Hubble parameter, $V_{b}$ is the potential associated with the homogeneous background, and $\tilde{\rho}_{\rm b}$ is the comoving background density. The similarities between Eqs. (12) and (45) allows the same numerical algorithms to be applied to both cases, particularly if one further introduces the so-called ‘supercomoving time’, $d\tilde{t}=a^{-2}dt$ (Martel and Shapiro 1998).
The wave function $\psi$ can be expressed in polar coordinates:
$\psi=R+iI=\rho^{1/2}e^{iS},$ (6)
where $R=\Re({\psi})$ and $I=\Im(\psi)$ are the real and imaginary parts of the wave function, respectively, and $S$ is the phase. One can then derive physical quantities, such as mass density $\rho$, bulk velocity $\bm{v}$, ‘thermal’ velocity $\bm{w}$, and energy density $e$, as
$\rho$ $=\|\psi\|^{2},$ (7)
$\bm{v}$ $=\frac{\hbar}{m}\frac{R\bm{\nabla}I-I\bm{\nabla}R}{\|\psi\|^{2}}=\frac{\hbar}{m}\bm{\nabla}S,$ (8)
$\bm{w}$ $=\frac{\hbar}{m}\frac{R\bm{\nabla}R+I\bm{\nabla}I}{\|\psi\|^{2}}=\frac{\hbar}{2m}\frac{\bm{\nabla}\rho}{\rho},$ (9)
$e$ $=\frac{\hbar^{2}}{2m^{2}}\|\bm{\nabla}\psi\|^{2}=\frac{1}{2}\rho(v^{2}+w^{2}).$ (10)
Similarly, in comoving coordinates, we can define $\tilde{\psi}=\tilde{\rho}^{1/2}e^{i\tilde{S}}$, with $\tilde{\rho}=a^{3}\rho$ being the comoving mass density. The peculiar velocity is then given by
$\bm{v_{\rm pec}}=a^{-1}\frac{\hbar}{m}\bm{\tilde{\nabla}}\tilde{S}$. (11)
1.1.2 Fluid formulation
From the transformation in Eq. (6), Eq. (1) can be recast into the Hamilton–Jacobi–Madelung equations:
$\frac{m}{\hbar}\frac{\partial\rho}{\partial t}+\bm{\nabla}\cdot(\rho\bm{\nabla}S)$ $=0,$ (12)
$\frac{m}{\hbar}\frac{\partial S}{\partial t}+\frac{1}{2}\|\bm{\nabla}S\|^{2}+\frac{m^{2}}{\hbar^{2}}\left(Q+V\right)$ $=0,$ (13)
where $Q$ is referred to as the ‘quantum potential’:
$Q=-\frac{\hbar^{2}}{2m^{2}}\frac{{\nabla}^{2}\sqrt{\rho}}{\sqrt{\rho}}=-\frac{\hbar^{2}}{2m^{2}}\left[\frac{\nabla^{2}\rho}{2\rho}-\frac{\|\bm{\nabla}\rho\|^{2}}{4\rho^{2}}\right].$ (14)
By taking the gradient of Eq. (13) and defining the bulk velocity $\bm{v}$ from Eq. (8), one obtains the Madelung equations:
$\frac{\partial\rho}{\partial t}+\bm{\nabla}\cdot\left(\rho\bm{v}\right)$ $=0,$ (15)
$\frac{\partial\bm{v}}{\partial t}+\bm{v}\cdot\bm{\nabla}\bm{v}+\bm{\nabla}(Q+V)$ $=0.$ (16)
Note that $\bm{v}\propto\bm{\nabla}S$ is a gradient flow, so the flow vorticity $\bm{\nabla}\times\bm{v}$ vanishes everywhere except at the location of vortices, where the density drops to zero and both the phase and velocity become ill-defined (see Section 3.2).
Eq. (16) can be rewritten into a conservative form:
$\frac{\partial\rho\bm{v}}{\partial t}+\bm{\nabla}\cdot\left(\rho\bm{v}\otimes\bm{v}+\bm{\Sigma}\right)+\rho\bm{\nabla}V=0,$ (17)
where $\bm{\Sigma}$ is the ‘quantum stress tensor’, often also referred to as the ‘quantum pressure’:
$\Sigma_{ij}=\frac{\hbar^{2}}{4m^{2}}\left(\frac{1}{\rho}\frac{\partial\rho}{\partial x_{i}}\frac{\partial\rho}{\partial x_{j}}-\delta_{ij}\nabla^{2}\rho\right)\quad\text{or}\quad-\frac{\hbar^{2}}{4m^{2}}\left(\rho\frac{\partial^{2}\ln\rho}{\partial x_{i}\partial x_{j}}\right),$ (18)
with $\delta_{ij}$ denoting the Kronecker delta. The two forms of $\Sigma_{ij}$ are equivalent after taking the divergence. Eq. (17) is analogous to the Euler equations of hydrodynamics, but with a key distinction: the quantum pressure term introduces several unique features in FDM (see Section 1.2).
[…]
1.2 FDM features
This subsection highlights key FDM features: the suppression of small-scale structure (Section 1.2.1), soliton cores (Section 1.2.2), and density granulation (Section 1.2.3).
1.2.1 Suppression of small-scale structure
Quantum pressure in FDM counteracts gravity and introduces a Jeans scale, which is associated with the de Broglie wavelength of the ground state in a potential well. Structures larger than this scale grow similarly to CDM, whereas smaller structures are suppressed due to enhanced velocity arising from the uncertainty principle. The comoving Jeans wavenumber $k_{\rm J}$ remains constant during the radiation-dominated era and increases slowly as $k_{\rm J}\propto m_{22}^{1/2}a^{1/4}$ during the matter-dominated era (Hu et al. 2000). This slowly varying $k_{\rm J}$ leads to a sharp transition in the linear matter power spectrum at $k\sim k_{\rm J,eq}$, where $k_{\rm J,eq}=9\,m_{22}^{1/2}\,\rm Mpc^{-1}$ is the Jeans wavenumber at matter-radiation equality.
The suppression of the FDM power spectrum $P_{\rm FDM}(k,z)$ relative to CDM can be expressed as
$P_{\rm FDM}(k,z)=T_{\rm FDM}^{2}(k,z)P_{\rm CDM}(k,z),$ (19)
where $T_{\rm FDM}$ is the FDM transfer function and $z$ is the redshift. Hu et al. (2000) proposed the following analytical fitting formula for $T_{\rm FDM}$ in the matter-dominated era:
$T_{\rm FDM}\approx\frac{\cos x^{3}}{1+x^{8}},\;\;x=1.61\,m_{22}^{1/18}\frac{k}{k_{\rm J,eq}}.$ (20)
It is modeled as redshift-independent because the strong suppression at $k\sim k_{\rm J,eq}$ is primarily determined during the radiation-dominated era. Alternatively, $T_{\rm FDM}$ can be computed directly using the Boltzmann code axionCAMB (Hložek et al. 2015).
From Eq. (20), one can define the ‘half-mode’ wavenumber $k_{1/2}$, where $T_{\rm FDM}(k_{1/2})=1/2$, and the associated halo mass $M_{1/2}=4\pi(\pi/k_{1/2})^{3}\tilde{\rho}_{\rm b}/3$ as
$k_{1/2}\approx 5.1\,m_{22}^{4/9}\,\rm Mpc^{-1},\;\;M_{1/2}\approx 3.8\times 10^{10}\,m_{22}^{-4/3}\,\rm{M}_{\odot}.$ (21)
$M_{1/2}$ characterizes the mass scale below which the FDM halo mass function, $\left.{\rm d}n/{\rm d}M_{\rm h}\right\rvert_{\rm FDM}$, is significantly suppressed compared to CDM. Schive et al. (2016) provided the following fitting function:
$\left.\frac{{\rm d}n}{{\rm d}M_{\rm h}}\right\rvert_{\rm FDM}(M_{\rm h},z)=\left.\frac{{\rm d}n}{{\rm d}M_{\rm h}}\right\rvert_{{\rm CDM}}(M_{\rm h},z)\left[1+\left(\frac{M_{\rm h}}{0.42M_{1/2}}\right)^{-1.1}\right]^{-2.2}.$ (22)
Similar to Eq. (20), the suppression term $1+(M_{\rm h}/0.42M_{1/2})^{-1.1}]^{-2.2}$ is redshift-independent, as $k_{\rm J}>k_{\rm J,eq}$ throughout the matter-dominated era. Furthermore, although Eq. (22) was originally derived from collisionless NN-body simulations with FDM initial conditions (see Section 2.6), it has been confirmed to agree well with genuine FDM wave simulations solving Eqs. (45) (May and Springel 2022). This agreement demonstrates that the FDM halo mass function in cosmological simulations is governed primarily by the initial conditions, with dynamical effects of quantum pressure playing a subdominant role.
Figure 1: Projected density distributions from three cosmological simulations at $z=0$: (left) a genuine CDM simulation, (middle) a collisionless NN-body simulation from an FDM initial condition (IC) with $m_{22}=0.1$, and (right) a genuine FDM simulation solving the Schrödinger–Poisson equations and their fluid formulation using the same FDM initial condition as the middle panel. Low-mass halos in the latter two cases are significantly suppressed in a very consistent manner, suggesting that the initial conditions dominate over the dynamical effects of quantum pressure in determining the FDM halo mass function
Fig. 1 illustrates these features by comparing the density distributions at $z=0$ from three distinct simulations: collisionless NN-body simulations initialized with CDM and FDM ($m_{22}=0.1$) power spectra, and a genuine FDM simulation using a hybrid fluid–wave algorithm (see Section 2.3) with the same FDM initial condition. The simulation box size is $5.9\,\rm Mpc$. The genuine CDM simulation exhibits significantly more low-mass halos. By contrast, the halo distributions in the latter two simulations are similar, although only the genuine FDM simulation can resolve small-scale interference fringes and solitons (see Section 1.2.2 and Section 1.2.3). The two surviving massive halos have a mass of $M_{\rm h}=4.5\times 10^{11}\,\rm{M}_{\odot}$ and $9.1\times 10^{11}\,\rm{M}_{\odot}$, respectively, close to the half-mode mass $M_{1/2}\approx 8.2\times 10^{11}\,\rm{M}_{\odot}$. The NN-body and FDM simulations are performed using the codes GADGET-2 (Springel 2005) and GAMER (Schive et al. 2018), respectively.
The suppression of small-scale structure provides a powerful means to constrain $m_{22}$. Examples include the Lyman-alpha forest power spectrum (Iršič et al. 2017; Armengaud et al. 2017; Kobayashi et al. 2017; Leong et al. 2019; Nori et al. 2019; Rogers and Peiris 2021), the cosmic microwave background (Hložek et al. 2015, 2018), the subhalo mass function (Du et al. 2017a; Benito et al. 2020; Schutz 2020; Banik et al. 2021; Nadler et al. 2021, 2025), the halo abundance and the faint end of the luminosity function at high redshifts (Bozek et al. 2015; Schive et al. 2016; Corasaniti et al. 2017; Menci et al. 2017; Schive and Chiueh 2018; Ni et al. 2019; Winch et al. 2024; Sipple et al. 2025), the galaxy shear weak lensing power spectrum (Dentler et al. 2022), and the 21-cm signal during cosmic dawn (Hotinli et al. 2022).
1.2.2 Solitons
Figure 2: Density distribution of a $5.7\times 10^{10}\,\rm{M}_{\odot}$ FDM halo with $m_{22}=0.2$ at $z=0$. The halo exhibits ubiquitous, stochastically fluctuating granular structures, with a characteristic granule size determined by the local de Broglie wavelength. The brightest region in the inset highlights the central dense soliton, which has a mass of $5.2\times 10^{8}\,\rm{M}_{\odot}$. The soliton radius is comparable to that of the surrounding granules, suggesting thermal equilibrium
FDM halos feature a dense, stable soliton core surrounded by a Navarro-Frenk-White (NFW; Navarro et al. 1996) halo permeated by fluctuating density granules (Schive et al. 2014a, b; Marsh and Pop 2015; Mocz et al. 2017; Veltmaat et al. 2018). Fig. 2 displays such a soliton–halo system. The soliton, forming shortly after halo collapse, represents the ground-state solution of the halo potential. The soliton density profile can be well fitted by (Schive et al. 2014a; Marsh and Pop 2015)
$\rho_{\rm s}(r)=\frac{1.95\times 10^{7}m_{22}^{-2}(r_{\rm s}/\,\rm kpc)^{-4}}{[1+9.06\times 10^{-2}(r/r_{\rm s})^{2}]^{8}}\,\rm{M}_{\odot}\,\rm kpc^{-3},$ (23)
which is redshift-independent as the soliton is an object detached from cosmological expansion. See Fig. 12 for an illustration. Here, $r_{\rm s}$ is the soliton radius defined by $\rho_{\rm s}(r_{\rm s})=\rho_{\rm s}(0)/2$. The corresponding soliton mass $M_{\rm s}$ within $r_{\rm s}$ is
$M_{\rm s}=5.4\times 10^{7}m_{22}^{-2}\left(\frac{r_{\rm s}}{\,\rm kpc}\right)^{-1}\,\rm{M}_{\odot}.$ (24)
Analytical expressions for the enclosed mass and gravitational potential profiles can be found in Eq. (A1) of Chen et al. (2017) and Eq. (6) of Chiang et al. (2021), respectively. Note that Eqs. (2324) obey the scale transformation
$(r_{\rm s},\rho_{\rm s},M_{\rm s},E_{\rm s})\to(\alpha^{-1}r_{\rm s},\alpha^{4}\rho_{\rm s},\alpha M_{\rm s},\alpha^{3}E_{\rm s}),$ (25)
which follows directly from Eq. (3) with $\beta=1$, where $E_{\rm s}$ denotes the soliton energy. Accordingly, for a fixed FDM particle mass, all soliton solutions form a one-parameter family.
Soliton properties can be inferred from their host halos, known as the soliton–halo relation. Its exact form remains under debate, with proposed scalings including $M_{\rm s}\propto M_{\rm h}^{1/3}$ (Schive et al. 2014b; Du et al. 2017b; Veltmaat et al. 2018; Chavanis 2019) and $M_{\rm s}\propto M_{\rm h}^{5/9}$ (Mocz et al. 2017; Nori and Baldi 2021). This relation can be interpreted as a manifestation of thermal equilibrium between the soliton and its surrounding halo granules (Liao et al. 2025). The transition from soliton to halo occurs at $r\approx 3.3\text{--}3.5\,r_{\rm s}$ (Mocz et al. 2017; Chiang et al. 2021), beyond which the average radial density profile follows the NFW model and closely matches the result from a collisionless NN-body simulation using the same FDM initial condition (Liao et al. 2025, see also the middle and right panels in Fig. 1). In contrast to CDM, the concentration parameter of FDM halos decreases for $M_{\rm h}\lesssim M_{1/2}$, as a consequence of the delayed onset of halo formation (Laroche et al. 2022; Kawaii et al. 2024).
The flat core and mass excess associated with the central soliton is a distinct prediction of FDM, setting it apart from other dark matter scenarios and providing a stringent test of the model. For instance, one can constrain the soliton profile from the size and kinematic data of dwarf galaxies (Schive et al. 2014a; Calabrese and Spergel 2016; Chen et al. 2017; González-Morales et al. 2017; Burkert 2020; Safarzadeh and Spergel 2020; Hayashi et al. 2021; Zimmermann et al. 2025), ultra-diffuse galaxies (Wasserman et al. 2019; Montes et al. 2024), and the central molecular zone of the Milky Way (Li et al. 2020). The deep gravitational potential of a dense compact soliton can (i) accelerate gas accretion and thereby boost the growth of supermassive black holes in the early universe (Chiu et al. 2025), and (ii) produce a characteristic peak in galaxy rotation curves at $r\approx 1.96\,r_{\rm s}$ with a peak velocity $v_{\rm max}\approx 36\,m_{22}(M_{\rm s}/10^{8}\,\rm{M}_{\odot})\,\rm km\,\rm s^{-1}$ (Bernal et al. 2018; Bar et al. 2018, 2019; Crǎciun and Harko 2020; Chan and Fai Yeung 2021; Bar et al. 2022; Khelashvili et al. 2023). This additional potential can also counteract tidal forces, making FDM subhalos more resilient to tidal disruption compared to their CDM counterparts (Chan et al. 2025). On the other hand, if the tidal radius is comparable to $r_{\rm s}$, solitons become unstable and undergo runaway disruption (Du et al. 2018). Furthermore, solitons exhibit density oscillations and random walk due to wave interference (Veltmaat et al. 2018; Schive et al. 2020; Li et al. 2021; Chiang et al. 2021; Dutta Chowdhury et al. 2021; Zagorac et al. 2022; Chiueh and Hsu 2023), which may dynamically heat or even tidally disrupt nearby stellar systems (Marsh and Niemeyer 2019; Schive et al. 2020; Dutta Chowdhury et al. 2023; Teodori et al. 2026).
1.2.3 Density granulation
As shown in Fig. 2, another distinct feature of FDM halos is the presence of ubiquitous, stochastically fluctuating granular structures, arising from constructive and destructive interference. The characteristic granule size is set by the local de Broglie wavelength $\lambda_{\rm dB}$ (Veltmaat et al. 2018; Bar-Or et al. 2019; Dutta Chowdhury et al. 2021),
$d_{\rm gra}\approx 0.25\,\lambda_{\rm dB}\approx 0.3\,m_{22}^{-1}\left(\frac{\sigma_{\rm 1D}}{100\,\rm km\,\rm s^{-1}}\right)^{-1}\,\rm kpc,$ (26)
where $\lambda_{\rm dB}=h/m\sigma_{\rm 1D}$ and $\sigma_{\rm 1D}$ is the one-dimensional (bulk) velocity dispersion. The characteristic granule lifetime is estimated as
$T_{\rm gra}\approx\frac{d_{\rm gra}}{\sigma_{\rm 1D}}\approx 2.9\,m_{22}^{-1}\left(\frac{\sigma_{\rm 1D}}{100\,\rm km\,\rm s^{-1}}\right)^{-2}\,\rm Myr.$ (27)
The dynamical effects of these density fluctuations can be modeled by treating FDM granules as quasiparticles (Bar-Or et al. 2019; Church et al. 2019; El-Zant et al. 2020; Chavanis 2021; Dutta Chowdhury et al. 2021). Assuming a Maxwellian velocity distribution, the effective particle mass is
$M_{\rm gra}\approx\frac{\pi\rho d^{3}}{6},$ (28)
where $\rho$ denotes the local dark matter density.
The granule size remains approximately constant in the vicinity of the soliton but increases with radius beyond this region, indicating an isothermal velocity distribution in the inner halo and a non-isothermal distribution in the outskirts. This radial variation in temperature correlates positively with the FDM halo concentration parameter (Liao et al. 2025). Furthermore, in contrast to the central soliton, which is supported by quantum pressure with $w\gg v$, the region outside the soliton exhibits $w\sim v$, indicating energy equipartition (Dutta Chowdhury et al. 2021). Since $\bm{w}\propto\bm{\nabla}\rho/\rho$ (see Eq. 9), the ubiquitous, large-amplitude density fluctuations can be interpreted as a manifestation of isotropic thermal velocity.
The perturbed gravitational field induced by FDM density granulation can scatter stars and gas, leaving observable imprints that may help distinguish FDM from other dark matter models. Examples include the outward diffusion of nuclear objects (Dutta Chowdhury et al. 2021) and stellar streams (Dalal et al. 2021), as well as the dynamical heating of star clusters (Marsh and Niemeyer 2019), dwarf galaxies (Dalal and Kravtsov 2022; Dutta Chowdhury et al. 2023; Teodori et al. 2026; Yang et al. 2025b, c), and galactic disks (Church et al. 2019; Chiang et al. 2023; Yang et al. 2024). These interference effects can also influence dynamical friction (Lancaster et al. 2020; Wang and Easther 2022; Vicente and Cardoso 2022; Vitsos and Gourgouliatos 2023; Foote et al. 2023), gravitational lensing (Chan et al. 2020; Laroche et al. 2022; Amruth et al. 2023; Powell et al. 2023), and pulsar timing signals (Khmelnitsky and Rubakov 2014; De Martino et al. 2017; Blas et al. 2017; Smarra et al. 2023). See also Hui et al. (2017); Bar-Or et al. (2019); El-Zant et al. (2020); Chavanis (2021) for further discussions.
2 Numerical methods
This section introduces a variety of FDM algorithms: wave-based schemes (Section 2.1), fluid-based schemes (Section 2.2), hybrid approaches (Section 2.3), adaptive mesh refinement (Section 2.4), eigenmode methods (Section 2.5), and collisionless NN-body methods (Section 2.6).
2.1 Wave-based methods
The time integration of the Schrödinger–Poisson equations can be approximated to second-order accuracy using a split-step method (Taha and Ablowitz 1984; Woo and Chiueh 2009; Mocz et al. 2017; Edwards et al. 2018; Angulo and Hahn 2022):
$\psi(\bm{r},t+\Delta t)$ $=\mathcal{T}\exp\left[-i\int^{t+\Delta t}_{t}{\rm d}t^{\prime}\left(-\frac{\hbar}{2m}\nabla^{2}+\frac{m}{\hbar}V(\bm{r},t^{\prime})\right)\right]\psi(\bm{r},t)$ (29)
$\approx\mathcal{K}(t+\Delta t,\Delta t/2)\,\mathcal{D}(t+\Delta t/2,\Delta t)\,\mathcal{K}(t,\Delta t/2)\,\psi(\bm{r},t)+\mathcal{O}(\Delta t^{3}),$ (30)
where $\mathcal{T}$ is the time ordering operator, and
$\mathcal{K}(t,\Delta t)$ $=\exp\left[-i\frac{m}{\hbar}\Delta tV(t)\right],$ (31)
$\mathcal{D}(t,\Delta t)$ $=\exp\left[i\frac{\hbar}{2m}\Delta t\nabla^{2}\right],$ (32)
represent the ‘kick’ and ‘drift’ operators, respectively. This method separates the nonlinear evolution due to the potential term (kick) from the linear evolution due to the kinetic term (drift). For a given wave function and potential at time $t$, the evolution proceeds by first applying the kick operator for a half time step, followed by the drift operator for a full time step. The potential is then recomputed at $t+\Delta t$, and a final half-step kick is applied. This procedure is analogous to the ‘kick-drift-kick (KDK)’ scheme commonly employed in NN-body simulations. To enhance computational efficiency, the initial kick of the current step and the last kick of the previous step can be combined into a single full-step update. Higher-order schemes are also available (Levkov et al. 2018; Schwabe et al. 2020).
The kick operator involves only algebraic operations in position space. Specifically, it rotates the phase of the wave function while leaving the density unchanged. The corresponding time-step constraint is commonly set by limiting the phase change per time step to below $2\pi$ (Feit et al. 1982), thereby avoiding phase aliasing:
$\Delta t_{\mathcal{K}}=\eta_{\mathcal{K}}\frac{2\pi\hbar}{m}\frac{1}{\|V\|_{\rm max}},$ (33)
where $\eta_{\mathcal{K}}\leq 1$ is a safety factor, and $\|V\|_{\rm max}$ denotes the maximum absolute value of the potential. However, since only phase differences are dynamically relevant, it may be more physically motivated to require that the relative phase variation between adjacent simulation elements per time step remain below $\pi$.
Similarly, the drift operator involves only algebraic operations in Fourier space, $\exp[-i\hbar\Delta tk^{2}/2m]$, where $k$ is the wavenumber. Requiring the phase change in the exponent to stay below $2\pi$ over a single time step yields the following time-step criterion:
$\Delta t_{\mathcal{D}}=\eta_{\mathcal{D}}\frac{4m}{\pi\hbar}\Delta h^{2},$ (34)
where the maximum wavenumber in one dimension is taken as $k_{\rm max}=\pi/\Delta h$. The safety factor satisfies $\eta_{\mathcal{D}}\leq 1$ in general, though its optimal value depends on the adopted wave scheme and specific application. The scaling $\Delta t_{\mathcal{D}}\propto\Delta h^{2}$ is characteristic of diffusion-like equations. From the fluid interpretation, $k_{\rm max}$ corresponds to the maximum bulk velocity resolvable at spatial resolution $\Delta h$ (see Eq. 8). Accordingly, this time-step condition is analogous to ensuring that a fluid element with the maximum velocity travels no more than one grid cell per time step.
In the following subsections, we introduce various wave schemes for solving the drift operator.
2.1.1 Global Fourier method
As mentioned above, the linear kinetic term in the Schrödinger equation becomes a scalar function in momentum space. Therefore, a pseudo-spectral method is arguably the algorithm of choice for integrating the drift operator. Specifically, it proceeds by first applying a discrete Fourier transform to the wave function, multiplying it by the drift operator in Fourier space, and then applying an inverse discrete Fourier transform to return the updated wave function to position space. This procedure can be schematically expressed as
$\mathcal{D}(t,\Delta t)=\mathcal{F}^{-1}\exp\left[-i\frac{\hbar}{2m}\Delta tk^{2}\right]\mathcal{F},$ (35)
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the discrete Fourier transform operator and its inverse, respectively. We refer to this approach as the ‘global Fourier method’ throughout this article to distinguish it from the local pseudo-spectral method introduced in Section 2.1.3. Representative codes implementing this method include GAMER (Schive et al. 2014a; Kunkel et al. 2025), AREPO (also referred to as AxiREPO; Mocz et al. 2017; May and Springel 2021), PyUltraLight (Edwards et al. 2018) and SPoS built on top of Enzo (Li et al. 2019).
The global Fourier method offers several distinct advantages. The use of the fast Fourier transform (FFT) makes it highly efficient on Central Processing Units (CPUs) and well suited for Graphics Processing Unit (GPU) acceleration. It achieves spectral convergence, with errors decaying faster than any algebraic rate as the number of grid points increases. Consequently, compared to other schemes, the global Fourier method can achieve comparable accuracy using significantly lower spatial resolution and larger time steps, particularly because $\Delta t\propto\Delta h^{2}$. Furthermore, the method is unconditionally stable (Taha and Ablowitz 1984), which, in principle, allows for using $\eta_{\mathcal{D}}>1$ in Eq. (34), depending on the accuracy requirements of the application. Finally, the method conserves mass to machine precision.
Despite these promising properties, the global Fourier method has a key limitation. The discrete Fourier transform assumes uniform spatial discretization and periodic boundary conditions. These constraints make the method unsuitable for use at refinement levels in adaptive mesh refinement (AMR) simulations, where compact stencils and Dirichlet or Neumann boundary conditions are typically required on locally refined grids. This limitation motivates the development of alternative approaches, such as finite-difference and local pseudo-spectral methods discussed in the following subsections.
2.1.2 Finite-difference methods
Finite-difference methods are well-suited for refined regions in AMR simulations and naturally support aperiodic boundary conditions. For example, the AxioNyx code (Schwabe et al. 2020) employs a fourth-order finite-difference method. The GAMER (Schive et al. 2014a) and SCALAR (Mina et al. 2020) codes expand the drift operator in position space as
$\mathcal{D}(t,\Delta t)=\sum_{p=0}^{q}\frac{1}{p\,!}\left[i\frac{\hbar}{2m}\Delta t\nabla^{2}\right]^{p}.$ (36)
The von Neumann stability analysis (Press et al. 2007) shows that this scheme is conditionally stable for $q\geq 3$. See also Section 2.2.2 for a finite-volume method implemented in the GIZMO code (Hopkins 2019).
The one-dimensional scheme along the $x$-direction with $q=3$ is expressed as
$\mathcal{D}(t,\Delta t)=1+i\left(\frac{\hbar}{2m}\Delta t\right)\frac{\partial^{2}}{\partial x^{2}}-\frac{1}{2}\left(\frac{\hbar}{2m}\Delta t\right)^{2}\frac{\partial^{4}}{\partial x^{4}}-\frac{i}{6}\left(\frac{\hbar}{2m}\Delta t\right)^{3}\frac{\partial^{6}}{\partial x^{6}}.$ (37)
The Laplacian operator can be approximated using a second-order finite-difference scheme:
$\frac{\partial^{2}}{\partial x^{2}}\psi^{n}_{j}=\frac{\psi^{n}_{j+1}-2\psi^{n}_{j}+\psi^{n}_{j-1}}{\Delta h^{2}},$ (38)
where $n$ and $j$ denote the time step and cell index, respectively. The resulting Courant–Friedrichs–Lewy (CFL) stability condition is
$\Delta t_{\mathcal{D}}=\frac{\sqrt{3}m}{2\hbar}\Delta h^{2},$ (39)
corresponding to $\eta_{\mathcal{D}}\approx 0.7$. Adopting a higher-order approximation for the Laplacian operator can reduce numerical errors, at the cost of larger stencils, increased computational cost, and a slightly more stringent CFL condition. Extension to three dimensions is straightforward using dimensional splitting, $\exp[\nabla^{2}]=\exp[\partial^{2}_{z}]\,\exp[\partial^{2}_{y}]\,\exp[\partial^{2}_{x}]$, where the order of operations is arbitrary since derivatives along different directions commute.
Finite-difference methods do not, in general, conserve mass, where the total mass is defined as $M_{\rm tot}=\sum_{j}\|\psi_{j}\|^{2}$. This issue can be resolved by additionally solving the continuity equation, Eq. (15). Here we describe a straightforward one-dimensional implementation as an illustration. First, the continuity equation can be discretized as
$\rho^{n+1}_{j}=\rho^{n}_{j}-\frac{\Delta t}{\Delta h}\left(F^{n+1/2}_{j+1/2}-F^{n+1/2}_{j-1/2}\right),$ (40)
where $F^{n+1/2}_{j\pm 1/2}$ are the mass fluxes at the right and left interfaces of cell $j$ at the half-step time $t+\Delta t/2$. To estimate the half-step fluxes, we note that the finite-difference scheme in Eq. (37) is equivalent to the following two-step formulation:
$\psi^{n+1/2}_{j}$ $=\left[1+\frac{i}{2}\left(\frac{\hbar}{2m}\Delta t\right)\frac{\partial^{2}}{\partial x^{2}}-\frac{1}{6}\left(\frac{\hbar}{2m}\Delta t\right)^{2}\frac{\partial^{4}}{\partial x^{4}}\right]\psi^{n}_{j},$ (41)
$\psi^{n+1}_{j}$ $=\psi^{n}_{j}+\left[i\left(\frac{\hbar}{2m}\Delta t\right)\frac{\partial^{2}}{\partial x^{2}}\right]\psi^{n+1/2}_{j}.$ (42)
Accordingly, we use the intermediate wave function $\psi^{n+1/2}_{j}$ from Eq. (41) to compute $F^{n+1/2}_{j\pm 1/2}$ using Eqs. (78):
$F^{n+1/2}_{j+1/2}$ $=\frac{\hbar}{m}\left(R^{n+1/2}_{j+1/2}\frac{\partial I^{n+1/2}_{j+1/2}}{\partial x}-I^{n+1/2}_{j+1/2}\frac{\partial R^{n+1/2}_{j+1/2}}{\partial x}\right),$ (43)
$R^{n+1/2}_{j+1/2}$ $=\frac{R^{n+1/2}_{j+1}+R^{n+1/2}_{j}}{2},$ (44)
$\frac{\partial R^{n+1/2}_{j+1/2}}{\partial x}$ $=\frac{R^{n+1/2}_{j+1}-R^{n+1/2}_{j}}{\Delta h},$ (45)
where $I^{n+1/2}_{j+1/2}$ and $\partial I^{n+1/2}_{j+1/2}/\partial x$ are defined analogously. Finally, we rescale the full-step solution $\psi^{n+1}_{j}$ from Eq. (42) to match the mass-conserving density $\rho^{n+1}_{j}$ from Eq. (40):
$\psi^{n+1}_{j,\rm rescaled}=\psi^{n+1}_{j}\frac{\sqrt{\rho^{n+1}_{j}}}{\|\psi^{n+1}_{j}\|}.$ (46)
Importantly, this fluid-based correction remains valid even in the presence of vortices, since the mass flux (Eq. 43) is well-behaved near nodes (see Section 3.2). Moreover, for AMR simulations, $F^{n+1/2}_{j\pm 1/2}$ can be reused to correct inter-level fluxes at resolution boundaries to preserve mass conservation. This capability is implemented in both GAMER and SCALAR. One caveat, however, is that this hybrid approach may degrade the error convergence rate and alter the stability condition of the original finite-difference methods, particularly when higher-order Laplacian approximations are employed.
The major drawback of finite-difference methods lies in its significantly slower error convergence compared to the global Fourier method (e.g., see Fig. 9). As a result, in AMR simulations employing the global Fourier method at the root level and a finite-difference method at refinement levels, it is typically necessary to activate multiple refinement levels simultaneously within a given region to surpass the accuracy of a root-level-only simulation (see Section 2.4 for further discussion). This limitation motivates the development of local pseudo-spectral methods, which use finite stencils like finite-difference methods but achieve comparable accuracy at lower resolution, as discussed in the next subsection.
2.1.3 Local pseudo-spectral methods
The global Fourier method is limited to periodic domains, while finite-difference methods exhibit much slower convergence. In addition, AMR simulations favor compact stencils to minimize computational and communication overhead. To meet these requirements, a variety of algorithms have been proposed to achieve fast convergence on non-periodic uniform grids. Here, we briefly introduce a local pseudo-spectral method based on Fourier continuations with Gram polynomials (Lyon 2009; Lyon and Bruno 2010), hereafter referred to as the ‘FC–Gram’ method. This method has been implemented in GAMER (Kunkel et al. 2025). It provides an efficient, accurate, and stable approach to extend non-periodic data onto a periodic grid, thereby enabling the use of a pseudo-spectral Fourier method and achieving high-order algebraic convergence.
Figure 3: Overview of the local pseudo-spectral FC–Gram method. Solid and dashed lines represent the wave function in the physical and periodically extended regions, respectively. The method consists of four main steps. (a) Filling ghost cells at the domain boundaries. (b) Extending the non-periodic data onto a periodic grid using Fourier continuations with Gram polynomials. (c) Evolving the wave function in the periodic domain using the discrete Fourier transform (DFT) with a spectral filter, where $C=-i\hbar\Delta t/2m$. (d) Discarding data in the ghost cells and the extended region. See text for details. Image reproduced with permission from Kunkel et al. (2025), copyright by AAS
Fig. 3 provides an overview of the FC–Gram method in one dimension, which consists of four main steps:
  1. (a)
    Filling ghost cells at the domain boundaries.
  2. (b)
    Extending the non-periodic data onto a periodic grid.
  3. (c)
    Evolving the wave function in the periodic domain.
  4. (d)
    Discarding data in the ghost cells and the extended region.
In Step (a), ghost-cell data of size $N_{\rm ghost}$ are added to each side of a target grid of size NN. These ghost cells can be filled by applying physical boundary conditions, copying from adjacent grids at the same resolution, or, in AMR simulations, interpolating from nearby coarse grids. This yields a physical domain of size $N+2N_{\rm ghost}$. Importantly, ghost cells serve a dual purpose. In addition to enabling data exchange among neighboring grids, they also help prevent unphysical data in the periodically extended region from contaminating the target grid, provided that an appropriate time-step constraint is imposed.
Step (b) is the most critical. It introduces smooth, periodic continuations of the non-periodic data to avoid the Gibbs phenomenon and enable the use of rapidly convergent Fourier methods. Specifically, it extends the non-periodic wave function from the physical domain into an additional region of size $N_{\rm ext}$, thereby forming a periodic computational domain of size $N_{\rm total}=N+2N_{\rm ghost}+N_{\rm ext}$. The wave function in this extended region is unphysical in the sense that it does not represent real simulation data. Instead, it is artificially constructed solely to create a smooth periodic connection between the two ends of the physical domain.
To generate this extension, the method first projects the wave function from $N_{\rm boundary}$ cells on each side of the physical domain onto Gram polynomials. These discrete orthogonal polynomials are then replaced with a set of Fourier modes that both match the Gram polynomials within the physical domain and remain periodic over the full $N_{\rm total}$ domain. The Fourier coefficients are determined by solving a linear optimization problem using singular value decomposition, which minimizes the mismatch between the Gram polynomials and their Fourier continuations within the physical domain. See Kunkel et al. (2025) for details.
In Step (c), the wave function in the full computational domain of size $N_{\rm total}$ is updated by a time step $\Delta t$. Since this extended domain is periodic, the wave function can be evolved straightforwardly using a pseudo-spectral Fourier method, similar to the scheme described in Section 2.1.1. There is, however, an important caveat: the artificial periodic extension can introduce high-$k$ modes that lead to numerical instability and pollute the wave function in the physical domain. This issue can be mitigated by truncating the Taylor expansion of the drift operator in Fourier space, $\exp[-i\hbar\Delta tk^{2}/2m]$, to order $N_{\rm Taylor}$ and further applying an exponential filter:
$\exp\left(-\mu(k/k_{\rm max})^{2\nu}\right)\sum_{p=0}^{N_{\rm Taylor}}\frac{(-i\hbar\Delta tk^{2}/2m)^{p}}{p!},$ (47)
where the filter parameters are set to $\mu=16\ln(10)$, $\nu=50$, and $k_{\rm max}=\pi/\Delta h$ by default (Albin and Bruno 2011). This combination of Taylor truncation and exponential filtering ensures numerical stability by suppressing spurious high-$k$ modes.
Finally, in Step (d), only the data within the original grid of size NN are retained.
The error convergence rate of the FC–Gram method depends on the order of the Gram polynomials. With the default configuration in GAMER, the method achieves eleventh-order accuracy. It thus significantly outperforms finite-difference methods, although it still falls short of the global Fourier method (see Fig. 9). Mass conservation, however, is not guaranteed to machine precision. This is because the pseudo-spectral Fourier update in Step (c) is unitary only on the full computational domain, with no guarantee of a discrete local conservation law. As a result, the effective inter-grid mass fluxes computed by adjacent grids generally do not match exactly. The empirically determined CFL condition is $\eta_{\mathcal{D}}\approx 0.2\text{--}0.3$. This time-step criterion must ensure not only numerical stability but also that the unphysical wave function in the periodically extended region does not propagate into the physical domain beyond the ghost boundaries. Finally, note that Steps (b)–(d) are all linear operations and can be consolidated into a single matrix multiplication, thereby improving computational efficiency for small grids.
2.2 Fluid-based methods
A key limitation in wave-based methods is the need to resolve the de Broglie wavelength even where the density is smooth and nonzero (see Section 3.1). Moreover, these methods do not in general guarantee conservation of mass, momentum, and energy, although mass conservation can be restored by additionally evolving the continuity equation. These limitations motivate fluid-based approaches, solving either the Hamilton–Jacobi–Madelung equations, Eqs. (1213), or the Madelung equations, Eqs. (1517). Fluid methods have several notable advantages over wave approaches. They require much lower resolution in regions with smooth density and velocity, enabling significantly larger simulation volumes. They ensure manifest conservation when solving the fluid equations in conservative forms. They can readily incorporate a Lagrangian formulation, thereby respecting Galilean invariance and supporting adaptive resolution. However, fluid approaches have inherent limitations—they cannot accurately resolve strong interference, since the velocity and quantum potential diverge at density nodes (see Section 3.2).
In the following, we introduce two families of fluid-based schemes: smoothed particle hydrodynamics (Section 2.2.1) and finite-volume/finite-difference methods (Section 2.2.2).
2.2.1 Smoothed particle hydrodynamics
Smoothed particle hydrodynamics (SPH) is widely used in astrophysical simulations (Monaghan 1992; Springel 2010). In SPH, the fluid is represented by Lagrangian particles. The physical quantity $F$ at the position of particle $a$, $F_{a}=F(\bm{r}_{a})$, is approximated as a kernel-weighted sum over neighbors: $F_{a}=\sum_{b}(m_{b}/\rho_{b})F_{b}W_{ab}(h)$, where $m_{b}$ and $\rho_{b}$ are the mass and density of neighbor $b$, and $W_{ab}(h)=W(\|\bm{r}_{a}-\bm{r}_{b}\|,h)$ is a smooth, spherically symmetric kernel function. Here, $h$ is the kernel (smoothing) length, which generally varies in space and time to keep the number of neighbors roughly constant. Given the particle-sampled fluid fields and their derivatives, the system is advanced in time by solving the hydrodynamic equations in Lagrangian form.
SPH appears readily applicable to FDM simulations by replacing the isotropic gas-pressure term with the quantum acceleration $\bm{\nabla}Q$ in Eq. (16). Accurately evaluating $\bm{\nabla}Q$ is, however, highly nontrivial because it involves third derivatives of the density field, which are prone to noise and numerical instability. Nori and Baldi (2018) proposed the following formulation implemented in the AX-GADGET code built on top of P-GADGET3:
$\frac{{\rm d}^{2}\bm{r}_{a}}{{\rm d}t^{2}}=-\bm{\nabla}Q_{a}=\frac{\hbar^{2}}{2m^{2}}\sum_{b}\frac{m_{b}}{f_{b}\rho_{b}}\left[\frac{\nabla^{2}\rho_{b}}{2\rho_{b}}-\frac{\|\bm{\nabla}\rho_{b}\|^{2}}{4\rho_{b}^{2}}\right]\bm{\nabla}W_{ab},$ (48)
where
$f_{b}=1+\frac{h_{b}}{3\rho_{b}}\frac{\partial\rho_{b}}{\partial h_{b}}$ (49)
is a correction accounting for variable smoothing lengths and ${\rm d}/{\rm d}t$ is the convective derivative. Eq. (48) follows from applying the variational principle to the discretized SPH Lagrangian, with Lagrange multipliers enforcing the constraint $\rho_{b}h_{b}^{3}={\rm const}$. The gradient and Laplacian of the density field are computed as
$\bm{\nabla}\rho_{a}$ $=\sum_{b}(\rho_{b}-\rho_{a})\frac{m_{b}\bm{\nabla}W_{ab}}{\sqrt{\rho_{a}\rho_{b}}},$ (50)
$\nabla^{2}\rho_{a}$ $=\sum_{b}(\rho_{b}-\rho_{a})\frac{m_{b}\nabla^{2}W_{ab}}{\sqrt{\rho_{a}\rho_{b}}}-\frac{\|\bm{\nabla}\rho_{a}\|^{2}}{\rho_{a}}.$ (51)
For other SPH-like formulations, see Mocz and Succi (2015); Veltmaat and Niemeyer (2016); Zhang et al. (2018b).
In principle, in addition to the conventional hydrodynamic time-step constraints, a diffusion-like criterion, $\Delta t\propto h^{2}$ (similar to Eq. 34), is required for numerical stability and accuracy. However, this condition has not been enforced in the aforementioned SPH codes; its impact thus remains to be assessed.
SPH offers several benefits owing to its Lagrangian nature. The method supports smoothly adaptive resolution and can be easily coupled to NN-body gravity solvers. It is Galilean invariant, which minimizes advection errors relative to Eulerian methods. However, it faces several significant challenges. The quantum potential $Q$ depends on higher-order spatial derivatives of the density, which can be very noisy with standard SPH derivative estimators. Both $Q$ and velocity diverge at vortices where $\rho\rightarrow 0$ (see Section 3.2 and Fig. 14), which can cause substantial truncation errors, hinder numerical convergence, and force prohibitively small time steps. The smoothing length, generally scaling as $h\propto\rho^{-1/3}$ (for fixed particle mass), can exceed the local de Broglie wavelength $\lambda_{\rm dB}$ in low-density regions, making it difficult to resolve density granulation and the associated quantum pressure within halos.
These limitations likely contribute to the following discrepancies between different SPH implementations and wave-based methods. First, it remains uncertain whether SPH can robustly resolve solitons and recover the soliton–halo relation in FDM halos. For example, the flat cores found in the SPH spherical-collapse and cosmological simulations of Zhang et al. (2018b, Fig. 4), Nori and Baldi (2021, Fig. 4), and Nori et al. (2023, Fig. 7) do not exhibit the sharp soliton–halo transition commonly predicted in wave-based simulations (e.g., see Fig. 13) and observed in the SPH soliton-collision simulations of Veltmaat and Niemeyer (2016, Fig. 4). Relatedly, the velocity dispersion within the flat cores in Zhang et al. (2018b, Fig. 5) and Nori et al. (2023, Fig. 8) remains roughly constant. This is inconsistent with the soliton solution, which is supported by quantum pressure rather than velocity dispersion, as confirmed by wave-based simulations (e.g., Liao et al. 2025). Furthermore, Nori and Baldi (2018) found that the matter power spectrum is suppressed on small scales by the quantum pressure relative to NN-body simulations. This contrasts with the results of the SPH simulations by Veltmaat and Niemeyer (2016) and the wave-based global spectral simulations by May and Springel (2021), which reported enhanced small-scale power at low redshifts due to wave interference. These limitations and controversies highlight ongoing uncertainties and raise concerns about the reliability of SPH simulations in capturing the coarse-grained properties of FDM.
2.2.2 Finite-volume and finite-difference methods
Hopkins (2019) presented a family of Lagrangian, meshless finite-volume and finite-mass Godunov schemes implemented in the GIZMO code to solve the Madelung equations, Eqs. (15) and (17). The simulation volume is partitioned into moving, unstructured meshes defined by a weighted kernel, resembling a Voronoi tessellation but with smoothed mesh boundaries. Each mesh carries volume-averaged conserved quantities, such as mass and momentum density, and moves with an arbitrary velocity $v_{\rm mesh}$. The conserved quantities are updated using a finite-volume formulation by estimating fluxes across mesh boundaries. If one sets $v_{\rm mesh}$ equal to the local dark matter bulk velocity $v$, the inter-mesh mass fluxes vanish, and the mass of each mesh remains constant during evolution. It thus eliminates the need to solve the continuity equation. This approach is known as the meshless finite-mass method.
The code employs a second-order, matrix-based gradient estimator, which is essential for accurately computing the quantum stress tensor that involves higher-order derivatives of the mass density. Time integration is performed using an explicit leapfrog scheme with adaptive time steps. In addition to the conventional CFL conditions in hydrodynamics, a time-step stability criterion similar to Eq. (34) is employed, with the grid spacing $\Delta h$ replaced by the kernel length $h$. Gravity is incorporated via operator splitting. Inter-mesh fluxes are computed using an HLL-type Riemann solver in the rest frame of the mesh interface, with the left and right states reconstructed using a piecewise-constant method. This scheme conserves mass and momentum to machine precision. To further ensure energy conservation, the method introduces an auxiliary scalar field that tracks the energy associated with ‘unresolved’ quantum potential arising from numerical dissipation. This energy is then coupled back into the momentum equation as an additional isotropic pressure term.
The meshless finite-volume/mass formulation provides several notable advantages. It shares the key properties of Lagrangian methods, as in SPH. It conserves mass, momentum, and (optionally) energy, even with adaptive time steps. Moreover, it is significantly more stable and accurate than SPH methods, especially at ‘well-behaved’ nodes where the velocity remains finite while the density approaches zero. This improvement is partly because the formulation evolves conserved quantities and the quantum stress tensor, both of which do not diverge at nodes (see Fig. 14 for an illustration). However, this approach still faces a common limitation of all fluid-based methods: difficulty handling ‘divergent’ nodes, such as vortices, where the velocity becomes singular (see Section 3.2 for further discussion). In addition, similar to SPH, its Lagrangian nature tends to concentrate resolution in high-density regions, making it difficult to resolve halo granules when $\lambda_{\rm dB}<h$. Consequently, it remains unclear how this approach performs relative to other fluid- and wave-based methods in more realistic, complex settings—for instance, whether it can resolve fine-grained wave structure or preserve coarse-grained FDM features in cosmological simulations.
A similar finite-volume method can be applied to evolve the wave function by noting that the Schrödinger equation can also be written in a flux-conservative form:
$\frac{\partial R}{\partial t}+\bm{\nabla}\cdot\left(\frac{\hbar}{2m}\nabla I\right)=\frac{m}{\hbar}VI,$ (52)
$\frac{\partial I}{\partial t}+\bm{\nabla}\cdot\left(-\frac{\hbar}{2m}\nabla R\right)=-\frac{m}{\hbar}VR,$ (53)
with gravity treated as a source term. However, this approach does not guarantee conservation of mass, momentum, or energy, due to their nonlinear dependence on the wave function. Mass conservation can be recovered by additionally solving the continuity equation and rescaling the wave function’s amplitude while keeping its phase unchanged, similar to the approach outlined in Section 2.1.2.
Kunkel et al. (2025) proposed a finite-difference method to solve the Hamilton–Jacobi–Madelung equations, Eqs. (1213), on the fluid levels of a hybrid scheme (see Section 2.3). The continuity equation is solved using a monotonic upstream-centered scheme for conservation laws (MUSCL), combined with a piecewise-linear data reconstruction and the van Albada limiter (van Albada et al. 1997). For the phase equation, the advection term $\|\bm{\nabla}S\|^{2}$ is computed using the Sethian–Osher flux (Osher and Sethian 1988), and the quantum potential term $\nabla^{2}\sqrt{\rho}/\sqrt{\rho}$ is discretized using the second-order central finite-difference method. Time integration is performed using a strong stability-preserving, third-order Runge–Kutta method (Shu 2007). The time-step criteria depend on both Eq. (34) and the spatial derivative of the phase field. This method conserves mass but not momentum or energy.
2.3 Hybrid methods
Fluid-based schemes are significantly more efficient than wave-based schemes for addressing smooth, high-velocity flows, which occupy most of the simulation volume outside collapsed regions in cosmological simulations. Nevertheless, fluid schemes cannot robustly resolve strong interference in multi-stream regions, where wave schemes are preferred. This dilemma motivates hybrid approaches, which apply fluid schemes in smooth, single-stream regions and switch to wave schemes in interference-dominated, multi-stream regions. In addition, hybrid schemes facilitate cosmological zoom-in simulations, as purely wave-based schemes cannot accurately capture the large-scale structure when the de Broglie wavelength outside the zoom-in regions is poorly resolved. The Lagrangian zoom-in regions can be identified either by performing collisionless NN-body simulations or by introducing tracer particles into FDM simulations, analogous to Bohmian mechanics (Wyatt 2005).
One critical challenge in such hybrid approaches is handling the fluid–wave interfaces—specifically, how to define a unique and accurate mapping between fluid variables and the wave function to impose consistent boundary conditions on both sides. Consequently, in fluid regions, it is generally preferable to evolve the phase field via the Hamilton–Jacobi–Madelung equations, Eqs. (1213), rather than evolving the velocity field via the Madelung equations, Eqs. (1516), since converting velocity to wave function is non-unique and subject to a time-dependent phase factor. The initial phase field can be computed by solving Eq. (83) given the velocity distribution.
Veltmaat et al. (2018) implemented a hybrid NN-body–wave method in the AMR code Enzo. On coarser AMR levels (treated as fluid regions), the quantum pressure is ignored and the dynamics are approximated by a collisionless NN-body approach. On the finest level, the code evolves the wave function using a finite-difference method. To reconstruct the wave function from NN-body particles, a ‘classical wave function’ formulation is adopted, in which each particle represents a localized wave packet with a central phase $S$. The phase of particle $a$ is updated by solving Eq. (13) in Lagrangian form without the quantum potential $Q$:
$\frac{{\rm d}S_{a}}{{\rm d}t}=\frac{m}{2\hbar}v_{a}^{2}-\frac{m}{\hbar}V(\bm{r}_{a}),$ (54)
where ${\rm d}/{\rm d}t=\partial/\partial t+(\hbar/m)\bm{\nabla}S\cdot\bm{\nabla}$ is the convective derivative. The phase of the wave function is reconstructed by superposing wave packets carried by particles using an interpolation kernel. The amplitude is obtained by applying the same kernel to particle mass (thus SPH-like) rather than to wave packets, thereby ensuring mass conservation but omitting interference effects. This reconstruction serves two purposes: (i) providing initial conditions for the wave solver in a target Lagrangian volume prior to shell-crossing; (ii) supplying boundary conditions at the fluid–wave interfaces for subsequent wave evolution. Conversely, no phase extraction from the wave function is required, because Eq. (54) involves no spatial derivatives of $S$.
One key limitation in this classical wave approximation is its inability to capture interference fringes. As a result, the initial wave function (on the highest AMR level) must be reconstructed in single-stream regions, and the fluid–wave interfaces must lie well outside the halo virial radius. To eliminate this restriction, Schwabe and Niemeyer (2022) proposed reconstructing the full wave function, including both phase and amplitude, from wave packets via a WKB-like Gaussian beam method with fixed amplitudes:
$\psi(\bm{r},t)=\sum_{a}W(\|\bm{r}-\bm{r}_{a}\|,h)\exp{\left[i\left(S_{a}+\frac{m}{\hbar}\bm{v}_{a}\cdot(\bm{r}-\bm{r}_{a})\right)\right]},$ (55)
where $W$ is a Gaussian kernel with smoothing length $h$. This improved method captures statistically correct interference patterns, enabling wave function reconstruction even in multi-stream regions well within the halo virial radius. However, it introduces mass conservation errors, which are mitigated by rescaling the average density of the wave function to match the underlying total particle mass. In addition, a time-step constraint is imposed to limit the maximum phase change per step in Eq. (54) to less than $\pi$. This hybrid method is implemented in the AMR code AxioNyx.
Figure 4: An FDM cosmological simulation using the hybrid fluid–wave scheme of Kunkel et al. (2025). The wave solver is activated at redshift $z=3.83$, prior to the development of interference around $z=3.1$. The top two rows show the projected density, with white grids indicating regions evolved by the wave solver. The bottom two rows present the zoom-in views of the density and unwrapped phase slices centered on the first wave region. Image reproduced with permission from Kunkel et al. (2025), copyright by AAS
In these NN-body–wave methods, the particles’ classical trajectories neglect the effects of quantum potential, and the reconstructed wave function is only accurate on a statistical level. To improve upon these limitations, Kunkel et al. (2025) proposed a fully grid-based hybrid scheme in the AMR code GAMER. On coarser levels, a finite-difference fluid scheme solving the Hamilton–Jacobi–Madelung equations in Eulerian form is employed (see Section 2.2.2). It retains the quantum potential term and thus remains accurate until divergent velocity and phase discontinuities develop at density voids. The phase difference between adjacent cells can substantially exceed $2\pi$ in general, reflecting the fact that fluid-based schemes do not need to resolve the de Broglie wavelength. Finer levels evolve the wave function using the local pseudo-spectral FC–Gram method described in Section 2.1.3. See Fig. 4 for an illustration.
Reconstructing the wave function from the fluid fields is unambiguous and straightforward, as both the density/phase and wave function are defined on grids, thus eliminating the need for an interpolation kernel. However, unlike in NN-body–wave methods, the fluid scheme here also requires boundary conditions at the fluid–wave interfaces, since Eq. (13) involves spatial derivatives of both $S$ and $\rho$. A complication arises from the fact that the phase extracted from the wave function is only defined up to a multiple of $2\pi$. To resolve this ambiguity, the fine-grid phase in the wave region is unwrapped to align smoothly with the coarse-grid phase in the fluid region. This phase unwrapping implicitly assumes that the de Broglie wavelength is resolved by at least two grid points across the fluid–wave interfaces, which necessitates sufficiently fine resolution before switching to the wave formulation. While this may appear to increase computational cost, it is in fact a prerequisite for any wave-based schemes to avoid significant aliasing errors. See Section 2.4 for the corresponding grid refinement criteria.
2.4 Adaptive mesh refinement
Uniform-resolution methods are inefficient for cosmological simulations due to the large dynamic range involved. This motivates the application of AMR to FDM (Schive et al. 2014a; Schwabe et al. 2020; Mina et al. 2020; Kunkel et al. 2025), which allows spatial and temporal resolution to adjust dynamically based on prescribed refinement criteria. However, designing robust and efficient refinement criteria for FDM simulations is nontrivial. Conventional quasi-Lagrangian or super-Lagrangian refinement strategies are insufficient, since the de Broglie wavelength (and thus the characteristic granule size) is generally independent of density. Velocity-based criteria are problematic near vortices, where the velocity diverges.
One common strategy is to apply the Löhner error estimator (Löhner 1987), based on the ratio of the second to the first derivatives, to both the real and imaginary parts of the wave function. In the hybrid scheme of Kunkel et al. (2025), a spectral refinement criterion based on the decay rate of the polynomial coefficients in the FC–Gram method (see Section 2.1.3) is applied to the wave levels. On the fluid levels, refinement is determined by the quantum potential and the Laplacian of the phase to detect strong interference and phase discontinuities.
Data exchange between different refinement levels involves interpolating coarse-grid data to fine grids and averaging fine-grid data to coarse grids, often referred to as the ‘prolongation’ and ‘restriction’ operations, respectively. Applying these operations directly to the wave function does not ensure mass conservation. This issue can be addressed by also interpolating/averaging the density field and then rescaling the amplitude of the wave function while preserving its phase, similar to the density correction method described in Section 2.1.2. Furthermore, for smooth, high-velocity flows, commonly found outside halos, interpolating/averaging the smooth density and phase fields gives more accurate results than operating directly on the rapidly oscillating wave function. However, the phase is discontinuous near vortices, where phase interpolation fails. Accordingly, Kunkel et al. (2025) proposed a hybrid interpolation scheme that switches between density/phase and wave-function interpolation based on the local second derivative of the phase field. Conservation errors can also arise at coarse–fine resolution interfaces, which, when solving the conservative fluid formulation, can be corrected by reconciling mismatches between fine- and coarse-grid fluxes, as commonly applied in hydrodynamic AMR simulations.
AMR supports adaptive time steps: lower-resolution levels can take larger time steps to reduce computational cost, which is particularly useful in FDM simulations due to the $\Delta t\propto\Delta h^{2}$ scaling. However, it introduces an underappreciated issue in wave-based, self-gravitating simulations—the boundary values of gravitational potential used on leaf patches for the Poisson solver can be incorrect. The reason is that these boundary values are obtained by interpolating the coarse-grid potential. Under adaptive time steps, coarse and fine grids are generally asynchronous, which necessitates temporal interpolation. This, in turn, requires advancing the density on coarser (non-leaf) patches before leaf patches, which is problematic for wave-based methods because effective mass fluxes can be highly inaccurate when the de Broglie wavelength is under-resolved. Applying the density correction scheme introduced in Section 2.1.2 does not resolve this issue, as the mass fluxes—and thus the resulting coarse-grained mass distribution—remain inaccurate, even though the total mass is conserved. One possible improvement is to compute coarse-grid fluxes by averaging those computed on leaf patches, rather than deriving them from the coarse-grid wave function. Another potential solution is to adopt a momentum-conserving wave-based scheme, which, to our knowledge, has yet to be developed. Note that this issue does not arise in conventional hydrodynamic AMR simulations, for which finite-volume methods ensure manifest momentum conservation regardless of resolution.
2.5 Eigenmode methods
Fully self-consistent Schrödinger–Poisson solvers can be computationally prohibitive, especially for large $m_{22}$. Furthermore, when studying the properties and impacts of FDM interference in isolated, nearly static, dark-matter-dominated systems, it is often reasonable to neglect the gravitational backreaction from non-dark-matter components. These considerations motivate the use of eigenmode-based methods, which approximate the time-dependent wave function as a linear superposition of eigenmodes of a static, smooth gravitational potential $V_{0}(\bm{r})$, thereby enabling efficient modeling of FDM halos that incorporate interference substructures.
The method involves three main steps:
  1. (a)
    For a target smooth density distribution $\rho_{0}(\bm{r})$, compute the corresponding gravitational potential $V_{0}(\bm{r})$.
  2. (b)
    Compute a library of bound-state eigenmodes associated with $V_{0}(\bm{r})$.
  3. (c)
    Determine a suitable superposition of eigenmodes such that the time-averaged density distribution best approximates $\rho_{0}(\bm{r})$ and remains stable.
These steps can be iterated, using the output density from Step (c) as the updated input $\rho_{0}(\bm{r})$ for Step (a), until convergence is achieved.
The expansion in Step (c) reads
$\psi(\bm{r},t)=\sum_{j}a_{j}\psi_{j}(\bm{r})e^{-iE_{j}t/\hbar},$ (56)
where $\psi_{j}$ is the $j\text{-th}$ eigenmode with energy eigenvalue $E_{j}$, and $a_{j}$ is the corresponding time-independent complex amplitude with a random phase. For a spherically symmetric potential $V_{0}(r)$, one commonly adopts $\psi_{j}=R_{n\ell}(r)Y_{\ell}^{m}(\theta,\phi)$, where $R_{n\ell}$ is the radial function, $Y_{\ell}^{m}$ are the spherical harmonics, and $n,\ell,m$ are the radial, angular momentum, and magnetic quantum numbers, respectively. The corresponding energy eigenvalue $E_{j}=E_{n\ell}$ is independent of $m$ due to spherical symmetry. Crucially, rather than numerically solving Eqs. (12), the wave function can be evolved analytically by propagating the phase of each eigenmode via $\exp(-iE_{j}t/\hbar)$, offering significant computational savings.
The density field is given by
$\rho(\bm{r},t)$ $=\|\psi(\bm{r},t)\|^{2}$ (57)
$=\sum_{j}\|a_{j}\|^{2}\|\psi_{j}(\bm{r})\|^{2}+\sum_{j\neq k}a_{j}a_{k}^{*}\psi_{j}(\bm{r})\psi_{k}^{*}(\bm{r})e^{i(E_{k}-E_{j})t/\hbar}.$ (58)
The first term on the right-hand side represents the time-independent average profile, whereas the second term describes the time-dependent interference patterns (i.e., density granulation), which statistically average out over time due to the random phases. The task in Step (c) is therefore to determine the coefficients $a_{j}$ such that the sum $\sum_{j}\|a_{j}\|^{2}\|\psi_{j}(\bm{r})\|^{2}$ matches the target smooth density $\rho_{0}(\bm{r})$ as closely as possible for self-consistency. Below, we briefly review several proposed methods to achieve this.
Lin et al. (2018) adopted the fermionic King model (Chavanis 1998) to assign the coefficients $a_{j}$. This approach arises from comparing several theoretical distribution functions with $a_{j}$ directly measured from halos in FDM cosmological simulations. The discrepancy between the density profiles in Steps (a) and (c) is corrected iteratively by treating the potential difference as a perturbed Hamiltonian. The shape of the central core (i.e., the soliton) is determined by the ground-state solution, while its amplitude is set by the soliton–halo relation (Schive et al. 2014b). One limitation of this method is the inability to specify arbitrary density profiles, as the final converged profile may differ noticeably from the initial guess. Furthermore, since the adopted theoretical distribution function is calibrated using simulation halos in a limited halo mass range, $2.8\times 10^{9}\,\rm{M}_{\odot}\text{--}7.0\times 10^{10}\,\rm{M}_{\odot}$, and at a fixed FDM particle mass, $m_{22}=0.81$, it is questionable whether the method remains reliable beyond this range. These concerns motivate the development of the alternative methods discussed below.
Dalal et al. (2021) adopted the Widrow–Kaiser ansatz (Widrow and Kaiser 1993), which constructs the wave function as a superposition of plane waves weighted by a classical distribution function $f(\bm{r},\bm{p})$:
$\psi(\bm{r})=m^{1/2}\sum_{\bm{p}}\left[f(\bm{r},\bm{p})d\bm{p}\right]^{1/2}N_{\bm{p}}e^{i\bm{r}\cdot\bm{p}/\hbar}.$ (59)
Here, $d\bm{p}$ is the discrete momentum-space resolution, and $N_{\bm{p}}$ are complex random numbers satisfying $\left\langle{N_{\bm{p}}^{*}N_{\bm{q}}}\right\rangle=\delta_{\bm{pq}}$, where $\left\langle{}\right\rangle$ denotes the ensemble average over different realizations. The distribution function $f(\bm{r},\bm{p})$ must satisfy the self-consistency condition: $m\sum_{\bm{p}}f(\bm{r},\bm{p})d\bm{p}=\rho_{0}(\bm{r})$. Under the assumptions of spherical symmetry and isotropy, $f(\bm{r},\bm{p})$ can be derived from the Eddington formula (Binney and Tremaine 2008; Teodori et al. 2026). To evolve the wave function using Eq. (56), the coefficients $a_{j}$ can be obtained by projecting the wave function from Eq. (59) onto the eigenmodes $\psi_{j}$. Unlike the method of Lin et al. (2018), this approach supports more general density profiles. One caveat, however, is that $f(\bm{r},\bm{p})$ cannot accurately model the central soliton core. This limitation may be addressed by superposing a theoretical soliton wave function following the soliton–halo relation. Finally, we note that Eq. (59) can also be used to construct a uniform FDM background with density granulation by inserting a Maxwell-Boltzmann distribution into $f(\bm{r},\bm{p})$ (Lancaster et al. 2020).
Yavetz et al. (2022) generalized the classical, particle-based Schwarzschild method (Schwarzschild 1979) to construct FDM halos by replacing particle orbits with wave eigenmodes (see also Yang et al. 2025a; Zimmermann et al. 2025). The coefficients $a_{j}$ are determined by formulating an optimization problem that minimizes the difference between the output density from Step (c) and the input $\rho_{0}(\bm{r})$ in Step (a). Additional constraints can be imposed during the optimization—for example, requiring $a_{j}$ to depend only on energy, analogous to an isotropic distribution function in particle-based models. A random phase is assigned to each $a_{j}$, mirroring the random orbital phases in the original Schwarzschild method. In the halo outskirts, where the Wentzel–Kramers–Brillouin (WKB) approximation holds, this approach is consistent with the Widrow–Kaiser ansatz. In contrast, in the central region, where the WKB-limit breaks down and the distribution function derived from the Eddington formula becomes invalid, the Schwarzschild method remains applicable and can support rather arbitrary cored profiles, which need not resemble conventional soliton-like solutions.
Compared to Schrödinger–Poisson solvers, eigenmode methods are significantly more computationally efficient. They also allow for simulating a restricted region of the halo without concern for boundary effects. However, evolving the wave function via Eq. (56) is only approximate, as it neglects the nonlinear interactions between eigenmodes and the gravitational backreaction from external objects.
This class of methods has been applied to a variety of studies, including the temporal and spatial correlation of density granules (Lin et al. 2018), soliton random walk, oscillations, and distortions (Li et al. 2021; Zagorac et al. 2022), and granulation-induced dynamical heating of stellar streams (Dalal et al. 2021) and ultra-faint dwarf galaxies (Dalal and Kravtsov 2022). In addition, eigenmode methods can be used solely to generate stable FDM initial conditions for Schrödinger–Poisson solvers. Example applications include studies of dynamical friction (Lancaster et al. 2020), soliton–halo relation (Yavetz et al. 2022), and granular heating of galactic disks (Yang et al. 2024) and dwarf spheroidal galaxies (Teodori et al. 2026).
2.6 Collisionless NN-body methods
The conventional collisionless NN-body methods offer a convenient and efficient way to capture the large-scale dynamics of FDM, as the Schrödinger–Poisson equations reduce to the Vlasov–Poisson equations in the limit $\lambda_{\rm dB}\to 0$ (Widrow and Kaiser 1993). In practice, the simulation setup follows that of CDM simulations, except that the initial density power spectrum is generated using the FDM transfer function (Eqs. 19 and 20), making it analogous to warm dark matter (WDM) simulations. Compared to fully wave-based FDM simulations, this NN-body approach allows for much larger comoving volumes due to its numerical efficiency. However, by neglecting quantum pressure, it fails to capture fine-grained wave phenomena, such as density granulation and soliton cores.
The NN-body methods, though approximate, have a number of practical applications. They can probe the FDM halo mass function in large-scale structure (Schive et al. 2016; May and Springel 2022), as the suppression of low-mass FDM halos is mainly determined by the simulation initial conditions (see Fig. 1). They are also useful for studying the morphology of FDM halos (Dome et al. 2023). By comparing approximate NN-body and genuine FDM simulations with the same initial conditions, one can quantify the dynamical effects of quantum pressure (Veltmaat and Niemeyer 2016; Nori and Baldi 2018; May and Springel 2021) and assess the numerical convergence of genuine FDM simulations (Liao et al. 2025, see also Fig. 16). In hybrid fluid–wave methods, the NN-body approaches can provide the large-scale gravitational field and the boundary conditions for wave regions (see Section 2.3), and can also be used to identify the Lagrangian volumes of selected halos in cosmological zoom-in simulations (Veltmaat et al. 2018; Chan et al. 2025).
Several studies have attempted to constrain the FDM particle mass using NN-body simulations, such as through the Lyman-alpha forest power spectrum (e.g., Iršič et al. 2017) and the subhalo mass function (e.g., Nadler et al. 2025). However, it remains unclear whether omitting strong interference fringes and quantum pressure effects—such as their impact on the tidal evolution of subhalos and solitons (Chan et al. 2025)—may bias these constraints. Moreover, NN-body approaches are susceptible to spurious halo formation, albeit with properties differing quantitatively from those in genuine FDM simulations (see Fig. 8).
3 Numerical challenges
This section examines key numerical challenges in FDM simulations: the small time and length scales associated with the de Broglie wavelength (Section 3.1), the singular behavior at vortices (Section 3.2), and the formation of spurious halos (Section 3.3).
3.1 de Broglie wavelength
Figure 5: Slices of the density (left) and phase (right) fields through the center of a $M_{\rm h}=4.5\times 10^{11}\,\rm{M}_{\odot}$ halo at $z\sim 0.8$ with $m_{22}=0.1$. Outside the halo, the density remains smooth while the phase exhibits rapid oscillations, corresponding to plane-wave solutions of the wave function. This requires wave-based schemes to employ much higher resolution than fluid-based schemes in order to accurately resolve the phase structure. Inside the halo, strong interference leads to ubiquitous, large-amplitude density fluctuations, necessitating high resolution for both wave-based and fluid-based schemes to resolve the de Broglie wavelength
The necessity of resolving the short de Broglie wavelength and its rapid oscillations associated with high-velocity flows poses a significant challenge for FDM simulations, especially for large $m_{22}$. The manifestation of the de Broglie wavelength in FDM cosmological simulations can be broadly categorized into two regimes: (i) plane waves in low-density regions outside halos, and (ii) strong interference fringes within halos and along filaments. See Fig. 5 for an illustration.
The large-scale, high-velocity flows in the free-fall regions outside halos correspond to plane-wave solutions of the wave function, $\psi\propto\exp(i\bm{k}\cdot\bm{r}-\omega t)$, with a wavelength $\lambda_{\rm dB}=h/mv\approx 1.2\,m_{22}^{-1}(v/100\,\rm km\,\rm s^{-1})^{-1}\,\rm kpc$. As a result, unlike CDM simulations, which typically employ high resolution only in dense regions, FDM cosmological simulations require high resolution even in low-density regions, where the density is smooth but the velocity is high. Failure to resolve the de Broglie wavelength associated with these high-velocity flows can lead to an underestimation of the velocity field, resulting in delayed structure formation. These low-density regions occupy a substantial fraction of the simulation volume, which is the primary reason why wave-based FDM simulations reaching lower redshifts are typically limited to comoving box sizes of only a few Mpc. By contrast, FDM simulations using fluid-based schemes can largely overcome this limitation, since the density, velocity, and phase fields are all smooth outside halos, allowing coarser resolution without resolving the de Broglie wavelength.
Strong interference leads to widespread, isotropic density granulation within virialized halos, as well as long, thin fringes along filaments. In contrast to plane waves, which exhibit oscillating phase but smooth density, the density field in these interference regions also displays large-amplitude fluctuations on the scale of the de Broglie wavelength associated with the local velocity dispersion (see Eq. 26). As a result, both fluid-based and wave-based schemes require similarly high spatial resolution. For example, for a $M_{\rm h}\sim 10^{12}\,\rm{M}_{\odot}$ halo with $\sigma_{\rm 1D}\sim 100\,\rm km\,\rm s^{-1}$ and $m_{22}=10$, one would need a minimum spatial resolution of at least $\sim 10\,\rm pc$. Insufficient resolution can lead to inaccurate estimation of the quantum potential and quantum pressure, both of which involve the second spatial derivative of the density (see Eqs. 14 and 18), potentially resulting in unphysical halo contraction or expansion and inaccurate soliton properties (Liao et al. 2025). Furthermore, density voids generated by strong destructive interference are ubiquitous within halos and can lead to the formation of vortices. In these zero-density regions, both the velocity and quantum potential diverge, making fluid-based schemes particularly vulnerable to large numerical errors. See Section 3.2 for further discussion.
Virialized FDM halos are approximately isothermal. For example, for a halo of mass $\sim 10^{12}\,\rm{M}_{\odot}$, the temperature in the central dense region is only about a factor of 2 higher than that near the virial radius, despite a density contrast of several orders of magnitude (e.g., see Liao et al. 2025). This near-isothermality implies that consistently high spatial resolution is required throughout the entire halo, resulting in substantial computational cost. Moreover, this requirement for nearly uniform resolution poses a particular challenge for Lagrangian approaches, such as SPH and moving mesh methods, whose spatial resolution typically scales with local mass density as $\rho^{-1/3}$. As a result, these methods may lack sufficient resolution in the low-density outskirts of halos, making it difficult to resolve density granules and, in turn, the quantum pressure in those regions.
The requirement of high spatial resolution for resolving density granules also necessitates high temporal resolution to capture their rapid oscillations, as $T_{\rm gra}\propto m_{22}d_{\rm gra}^{2}$ (see Eq. 27). This constraint applies equally to both fluid and wave schemes.
The computational costs as functions of halo mass $M_{\rm h}$ and FDM particle mass $m_{22}$ can be estimated as follows. The characteristic velocity dispersion scales as $\sigma_{\rm 1D}\propto(M_{\rm h}/r_{\rm h})^{1/2}\propto M_{\rm h}^{1/3}$, where $r_{\rm h}\propto M_{\rm h}^{1/3}$ is the halo virial radius. The spatial and temporal resolution required to resolve density granulation, $\Delta h$ and $\Delta t$, thus scale as
$\Delta h$ $\propto(m_{22}\sigma_{\rm 1D})^{-1}\propto m_{22}^{-1}M_{\rm h}^{-1/3},$ (60)
$\Delta t$ $\propto m_{22}\Delta h^{2}\propto m_{22}^{-1}M_{\rm h}^{-2/3}.$ (61)
We assume here that the simulation time steps are dominated by the drift operator, Eq. (34), which generally holds within halos. The memory consumption, $C_{\rm mem}$, is proportional to the total number of simulation elements $N\propto(r_{\rm h}/\Delta h)^{3}$. The simulation time, $C_{\rm time}$, scales as $N/\Delta t$, assuming that each time step requires $\mathcal{O}(N)$ operations rather than, e.g., $\mathcal{O}(N\ln N)$. These yield
$C_{\rm mem}$ $\propto m_{22}^{3}M_{\rm h}^{2},$ (62)
$C_{\rm time}$ $\propto m_{22}^{4}M_{\rm h}^{8/3}.$ (63)
The strong dependence on $m_{22}$ and $M_{\rm h}$ implies that computational costs can quickly become computationally prohibitive for large $m_{22}$ and $M_{\rm h}$.
3.2 Vortices
Figure 6: An illustration of the numerical challenge in resolving vortices within regions of vanishing density. The two open circles represent the simulation elements. Although the infinitesimal wave function $\delta\psi$ varies smoothly near the vortex at the origin, the phase field $S$ exhibits rapid variation, and the corresponding velocity diverges. This steep phase gradient and divergent velocity pose significant challenges for fluid-based schemes
Strong interference in an FDM halo can generate zero-density regions. They appear as closed loops located at the intersection of the two isosurfaces $\Re({\psi})=0$ and $\Im(\psi)=0$, where $\Re({\psi})$ and $\Im(\psi)$ denote the real and imaginary parts of the wave function, respectively. Along these nodal lines, the phase field $S$ becomes multi-valued, rendering the velocity field $\bm{v}\propto\bm{\nabla}S$ ill-defined. As a result, the flow vorticity $\bm{\nabla}\times\bm{v}$ is a Dirac delta function, and the velocity circulation $\oint\bm{v}\cdot d\bm{l}$ enclosing such a nodal line is quantized due to phase winding. These closed loops of zero density can thus be interpreted as vortex rings (Chiueh et al. 2011; Hui et al. 2021). The number density of vortex rings is approximately one per de Broglie volume, implying that vortices are ubiquitously distributed throughout each halo.
In the vicinity of a vortex line, the wave function remains well-behaved, but the phase field exhibits rapid variation. See Fig. 6 (and also Fig. 10) for an illustration. The density and velocity fields scale as $\rho\sim r^{2}$ and $v\sim r^{-1}$, respectively, where $r$ is the distance to the vortex line. Consequently, both the velocity and the quantum potential ($Q\propto{\nabla}^{2}\sqrt{\rho}/\sqrt{\rho}$) diverge as $r\rightarrow 0$ (see Fig. 14).
Figure 7: Wave function near a corotating vortex pair described by Eq. (68) at $t=0$. The two panels on the left present slices of density and phase, with insets showing a close-up of the vortices, which feature zero density and discontinuous phase. The panels on the right display profiles along lines directly through (solid) and slightly below (dashed) the vortex pair (see insets on the left). While the density and wave function vary smoothly across a vortex, the phase field exhibits a $\pi$ jump exactly at the vortex and changes rapidly nearby (see also Fig. 6). This steep phase gradient cannot be alleviated by the singular gauge transformation, $\rho^{\prime}=-\rho$ and $S^{\prime}=S+\pi$. Image reproduced with permission from Kunkel et al. (2025), copyright by AAS
As an example, Fig. 7 shows the distribution of wave function, phase, and density near a vortex pair described by Eq. (68) at $t=0$. The wave-based schemes work well for vortices since both $\Re({\psi})$ and $\Im(\psi)$ vary smoothly. However, the discontinuity in the phase field is numerically problematic for simulations evolving the Hamilton–Jacobi–Madelung equations, Eqs. (1213). Although applying a singular gauge transformation, $\rho^{\prime}=-\rho$ and $S^{\prime}=S+\pi$, can regularize both the density and phase exactly at the vortex, the phase still varies rapidly in its vicinity. Increasing spatial resolution offers limited improvement, if any, because the phase gradient becomes even steeper closer to the vortex, as $\|\bm{\nabla}S\|\propto r^{-1}$. Similarly, the divergence of the velocity and quantum potential poses significant challenges for simulations based on the Madelung equations, Eqs. (1516). A further complication arises for Lagrangian approaches, whose inherently lower resolution in underdense regions makes it difficult to resolve vortices accurately.
Solving the momentum conservation equation, Eq. (17), may seem promising at first glance, since at vortices the momentum density $\rho v$ vanishes, and both the momentum density flux $\rho v^{2}$ and the stress tensor $\Sigma_{ij}\propto\rho\frac{\partial^{2}\ln\rho}{\partial x_{i}\partial x_{j}}$ remain finite (as shown in Fig. 14). However, computing the momentum density flux involves evaluating the ratio of two infinitesimal quantities, $(\rho v)^{2}$ and $\rho$. Likewise, an accurate estimation of the stress tensor depends on a delicate cancellation between the divergent term $\frac{\partial^{2}\ln\rho}{\partial x_{i}\partial x_{j}}$ and the vanishing term $\rho$. These calculations are therefore highly susceptible to large numerical errors.
In short, resolving vortices using fluid-based schemes or Lagrangian approaches requires the development of innovative algorithms. It remains unclear whether, and to what extent, the inability to capture local vortices would affect the intrinsic global properties of FDM halos, such as the granular halo structure and its associated dynamical heating effects, the amplitude of quantum pressure, the degree of energy equipartition between kinetic and thermal energies, and the soliton–halo relation.
3.3 Spurious halos
It is well established that WDM simulations suffer from the formation of low-mass spurious halos, resulting in an unphysical upturn at the low-mass end of the halo mass function (Wang and White 2007; Angulo et al. 2013; Schneider et al. 2013). These spurious halos arise from artificial fragmentation along filaments and are seeded by numerical artifacts. Consequently, their masses and positions are sensitive to numerical accuracy, and simulations with different resolutions generally yield distinct distributions of spurious halos. In addition, their protohalos at the initial redshift of a simulation tend to exhibit flattened, disk-like geometries. In comparison, the distribution of genuine halos shows good convergence with increasing simulation accuracy, and their protohalos are typically more spheroidal (Lovell et al. 2014).
Similar to WDM simulations, spurious halos with mass significantly below the half-mode mass, $M_{1/2}\propto m_{22}^{-4/3}$ (see Eq. 21), can also form in FDM simulations, especially at lower redshifts. One may wonder whether quantum pressure could alleviate this issue, given that it can suppress small-scale structure. However, the dynamical impact of quantum pressure during the simulation is largely confined to highly nonlinear regions with large-amplitude density fluctuations, such as within granular halos. This aligns with the fact that the suppression of low-mass FDM halos is mainly governed by the sharp cutoff in the linear density power spectrum established at matter-radiation equality (see Section 1.2.1). In filaments, velocity dispersion and quantum pressure are significant only in directions perpendicular to the filament axis and are negligible along it, as indicated by the interference fringes oriented parallel to filaments (see the bottom panel in Fig. 8). As a result, small-scale perturbations can become Jeans unstable along filaments, but these are primarily seeded by numerical noise rather than physical perturbations. This ultimately leads to the formation of spurious FDM halos.
Figure 8: An illustration of spurious halos in FDM simulations. The image compares the density distributions along a filament at $z=0$ across three different simulation setups from the same FDM initial condition: (top) a collisionless NN-body simulation, (middle) a genuine FDM simulation with AMR using a hybrid fluid–wave algorithm, and (bottom) a genuine FDM simulation using the global Fourier method. Spurious halos, highlighted by arrows in the top and middle panels, appear along the filament as regularly spaced clumps with masses substantially below the half-mode mass, $M_{1/2}\propto m_{22}^{-4/3}$. Notably, these halos lack consistent counterparts in simulations with different setups, as they arise from artificial fragmentation seeded by numerical noise. The bottom panel shows no evidence of spurious halos, presumably due to the superior accuracy of the global Fourier method
Fig. 8 illustrates this issue by comparing the density distributions along a filament across three different simulation setups, all starting from the same FDM initial condition with $m_{22}=0.8$: (i) a collisionless NN-body simulation (Section 2.6), (ii) a genuine FDM simulation using the hybrid fluid–wave algorithm based on the local pseudo-spectral FC–Gram method (Section 2.3), and (iii) a genuine FDM simulation using the global Fourier method (Section 2.1.1). The two massive halos at the ends of the filament are presumably genuine, as their masses are comparable to $M_{1/2}$ and they appear consistently in all three simulations. By contrast, the low-mass clumps along the filament in the upper and middle panels are likely artifacts, as they exhibit several features commonly associated with spurious halos. First, their masses are well below $M_{1/2}$. Second, their spatial distribution shows a suspiciously regular spacing along the filament. Third, these halos lack consistent counterparts in simulations with different setups, suggesting that they are sensitive to numerical accuracy.
Spurious halos in the NN-body simulation are expected, since this setup is closely analogous to a WDM simulation, differing only in the exact shape of the initial density power spectrum. Their emergence in the genuine FDM simulation (middle panel) is also not surprising, given the above discussion, although this issue has received comparatively less attention than in the WDM case. For example, do spurious protohalos in FDM simulations also exhibit flatter geometries than genuine halos? How do their characteristic mass and spacing depend on resolution and algorithm? Notably, Fig. 8 suggests that the global Fourier method may be free from artificial fragmentation, possibly due to its superior accuracy along smooth filaments. However, this method cannot capture fine structures within the two massive halos once the de Broglie wavelength approaches the grid scale.
4 Test bench
This section describes a set of representative numerical tests for both validating numerical accuracy and illustrating FDM features. These include Gaussian wave packets (Section 4.1.1), vortices (Section 4.1.2), the Jeans instability (Section 4.2.1), solitons (Section 4.2.2), isolated halos (Section 4.2.3), and cosmological simulations (Section 4.2.4). Other illustrative tests in the literature include oblique traveling waves (Hopkins 2019), the node problem in plane-wave interference (Hopkins 2019), simple harmonic oscillators (Mocz and Succi 2015), the self-similar solution of a density jump (Hui et al. 2017; Li et al. 2019), spherical collapse (Schwabe et al. 2020), and two-stream collisions (Hui et al. 2017).
4.1 Tests without self-gravity
4.1.1 Gaussian wave packets
The Gaussian wave packet provides a simple analytical solution to the free-particle Schrödinger equation. The wave function with mean momentum $p_{0}=mv_{0}$ and initial mean position $x_{0}$ is given by
$\psi(x,t)=\frac{A}{\sqrt{\Delta^{2}+\frac{i\hbar}{m}t}}\exp{\left[-\frac{(x-x_{0}-v_{0}t)^{2}}{2(\Delta^{2}+\frac{i\hbar}{m}t)}\right]}\exp{\left[i\frac{mv_{0}}{\hbar}\left(x-x_{0}-\frac{v_{0}}{2}t\right)\right]},$ (64)
where $\Delta$ represents the position uncertainty and $A=\Delta^{1/2}\pi^{-1/4}$ normalizes the total mass to unity. The corresponding mass density and bulk velocity, derived from Eqs. (78), are
$\rho(x,t)$ $=\frac{A^{2}}{\sqrt{\Delta^{2}\gamma(t)}}\exp{\left[-\frac{(x-x_{0}-v_{0}t)^{2}}{\gamma(t)}\right]},$ (65)
$v(x,t)$ $=v_{0}+\frac{(x-x_{0}-v_{0}t)\hbar^{2}}{\Delta^{2}m^{2}\gamma(t)}t,$ (66)
$\gamma(t)$ $=\Delta^{2}+\frac{\hbar^{2}}{\Delta^{2}m^{2}}t^{2}.$ (67)
Unlike a single plane wave, the Gaussian wave packet presents a nontrivial task for fluid-based schemes. It therefore serves as a useful benchmark for comparing wave and fluid schemes and for calibrating their respective optimal resolutions as functions of $v_{0}$ and $\Delta$. In particular, fluid schemes generally require much lower resolution when evolving a single wave packet with high $v_{0}$. However, in the case of two colliding wave packets, strong interference can lead to divergent velocity and quantum potential at nodes of zero density, posing numerical challenges for fluid schemes even at very high resolution (Li et al. 2019).
Figure 9: Test of a traveling Gaussian wave packet. (Left) Initial density and real part of the wave function. Crosses and circles represent simulation data from the wave and fluid schemes, with resolutions of $\Delta h\approx\lambda_{\rm dB}/3$ and $\lambda_{\rm dB}$, respectively. (Right) Simulation results at $t=0.2$ using different numerical schemes. The global Fourier method (circles) yields the most accurate result; the FC–Gram method (triangles) introduces mild distortion, while the wave-based finite-difference method (squares) shows significant deviation from the analytical solution (solid line). Notably, the fluid scheme (pentagons) successfully captures the overall wave packet motion, despite not resolving the de Broglie wavelength ($\Delta h\sim\lambda_{\rm dB}$)
Fig. 9 shows the evolution of a single traveling Gaussian wave packet with $\Delta=0.8$, $v_{0}=33$, $x_{0}=3.2$, $A=0.67$, and $m/\hbar=1$. The left panel displays the initial condition at $t=0$. The de Broglie wavelength associated with the bulk velocity, $\lambda_{\rm dB}=2\pi/v_{0}$, is much shorter than the Gaussian width, illustrating the advantage of fluid-based methods in this regime. The right panel compares numerical results at $t=0.2$ obtained using four different schemes: the global Fourier method, the local pseudo-spectral FC–Gram method, the wave-based finite-difference method, and the fluid-based finite-difference method solving the Hamilton–Jacobi–Madelung equation (see Section 2 for details). To test the robustness of each approach, a deliberately low resolution of approximately three cells per $\lambda_{\rm dB}$ is adopted for the wave methods. The fluid method, by contrast, uses a resolution three times lower and therefore does not resolve $\lambda_{\rm dB}$. No AMR is applied. As shown, the global Fourier method closely reproduces the analytical solution. The FC–Gram method introduces mild distortion, while the wave-based finite-difference method exhibits significant deviations. Notably, the fluid method performs well despite its substantially coarser resolution.
4.1.2 Vortices
As discussed in Section 3.2, vortices are ubiquitous in FDM halos, where the phase field is discontinuous and both the velocity and quantum potential diverge. As a result, vortices pose a serious challenge for fluid-based methods. In contrast, wave-based methods remain robust, as the wave function varies smoothly across vortices. In hybrid schemes, vortices can serve as benchmarks for calibrating the criteria of switching between fluid and wave methods and for monitoring potential numerical artifacts at the fluid–wave interfaces.
Figure 10: A corotating vortex pair described by Eq. (68). See text for details
As an example, Eq. (68) describes a corotating vortex pair:
$\psi(R,\varphi)=\rho_{0}-B{\rm J_{1}}\left(R\sqrt{\frac{2\omega m}{\hbar}}\right){\rm e}^{i(\varphi-\omega t)},$ (68)
where $\rm J_{1}$} is the Bessel function of the first kind, $\omega$ is the angular frequency, $\rho_{0}$ is the background density, $B$ is a constant, $R$ is the cylindrical radius, and $\varphi$ is the polar angle. Fig. 10 shows the analytical results for $\rho_{0}=2$, $B=5$, $\omega=90$, and $m/\hbar=1$ at $t=0.8$. Note that the $2\pi$ phase jump perpendicular to the line connecting the vortex pair can be uniquely unwrapped, provided the de Broglie wavelength is well resolved. However, the $\pi$ phase jump along this line poses numerical difficulties for fluid-based schemes (see also Fig. 7). See Hui et al. (2021) for a variety of vortex configurations.
4.2 Tests with self-gravity
4.2.1 Jeans instability
In analogy to the Jeans instability in hydrodynamics, quantum pressure in FDM gives rise to a characteristic Jeans scale. Perturbations on scales larger than this threshold grow under gravity, whereas those on smaller scales are stabilized by quantum pressure. To examine the linear evolution of a perturbed wave function in a homogeneous background, one can insert $\psi=(1+\delta R)+i\delta I$ into Eqs. (12) and assume $\delta R\ll 1$ and $\delta I\ll 1$, with the background density normalized to unity. By retaining the leading-order terms and applying the conventional Jeans swindle, the spatial Fourier components of the wave function satisfy
$\frac{{\rm d}^{2}\delta R_{k}}{{\rm d}t^{2}}$ $=\frac{\hbar^{2}}{4m^{2}}\left(k_{\rm J,phy}^{4}-k^{4}\right)\delta R_{k},$ (69)
$\delta I_{k}$ $=\frac{2m}{\hbar k^{2}}\frac{{\rm d}\delta R_{k}}{{\rm d}t},$ (70)
where $k$ is the wavenumber and
$k_{\rm J,phy}=\left(\frac{16\pi Gm^{2}}{\hbar^{2}}\right)^{1/4}$ (71)
is the Jeans wavenumber in physical coordinates (note that $k_{\rm J}$ refers to the comoving Jeans wavenumber throughout this article). The corresponding solutions are
$\delta R(x,t)$ $=\delta R_{0}\cos(kx+\theta_{0}){\rm e}^{\pm\omega_{1}t},\quad$ $k<k_{\rm J,phy},$ (72)
$\delta R(x,t)$ $=\delta R_{0}\cos(kx\pm\omega_{2}t+\theta_{0}),\quad$ $k>k_{\rm J,phy},$ (73)
where $\omega_{1}=(\hbar/2m)(k_{\rm J,phy}^{4}-k^{4})^{1/2}$, $\omega_{2}=(\hbar/2m)(k^{4}-k_{\rm J,phy}^{4})^{1/2}$, and $\delta R_{0}$ and $\theta_{0}$ are arbitrary constants.
Figure 11: Analytical solution $f(\xi)$ (see Eq. 77) to the comoving Jeans instability equation (Eq. 74), where $\xi\propto k^{2}a^{-1/2}$. The plot illustrates the transition from growth to oscillation at the comoving Jeans wavenumber $k_{\rm J}$, which corresponds to $\xi^{2}=6$ (vertical line; see Eq. 79)
Similarly, in comoving coordinates, one can insert $\tilde{\psi}=(1+\delta\tilde{R})+i\delta\tilde{I}$ into Eqs. (45) to obtain the following governing equations for the spatial Fourier components of the comoving wave function (Woo and Chiueh 2009):
$\xi^{2}\frac{{\rm d}^{2}\delta\tilde{R}_{k}}{{\rm d}\xi^{2}}$ $=(6-\xi^{2})\delta\tilde{R}_{k},$ (74)
$\delta\tilde{I}_{k}$ $=-\frac{{\rm d}\delta\tilde{R}_{k}}{{\rm d}\xi},$ (75)
where
$\xi=\frac{\hbar}{mH_{0}}\frac{k^{2}}{a^{1/2}}.$ (76)
$H_{0}$ denotes the present-day Hubble parameter. Here we assume a matter-dominated universe with matter density parameter $\Omega_{m}=1$ and normalize the background density $\tilde{\rho}_{\rm b}$ to unity. Eq. (74) admits two independent solutions:
$f(\xi)$ $=\frac{3\cos\xi+3\xi\sin\xi-\xi^{2}\cos\xi}{\xi^{2}},$ (77)
$g(\xi)$ $=\frac{3\sin\xi-3\xi\cos\xi-\xi^{2}\sin\xi}{\xi^{2}}.$ (78)
Furthermore, a characteristic length scale arises at $\xi^{2}=6$, which defines the comoving Jeans wavenumber (Hu et al. 2000):
$k_{\rm J}=\left(\frac{6aH_{0}^{2}m^{2}}{\hbar^{2}}\right)^{1/4}=a^{1/4}k_{\rm J,phy}.$ (79)
For $k\ll k_{\rm J}$, $f(\xi)\propto a$ and $g(\xi)\propto a^{-3/2}$, corresponding to the growing and decaying modes, respectively. For $k\gg k_{\rm J}$, both $f(\xi)$ and $g(\xi)$ exhibit oscillatory behavior. Fig. 11 illustrates $f(\xi)$ and the transition at $\xi^{2}=6$.
The Jeans instability setup is useful for quantifying the error convergence rates of both fluid-based and wave-based schemes involving self-gravity. The corresponding solutions for fluid variables can be derived straightforwardly from Eqs. (710) (see also Woo and Chiueh 2009). There is, however, one caveat. The solutions given in Eqs. (7778) rely on the assumption of $\delta\tilde{I}^{2}\ll 2\delta\tilde{R}$ in the source term of the Poisson equation. This assumption can break down rapidly in the growing-mode solution with $\xi\ll 1$, since $\|\delta\tilde{I}/\delta\tilde{R}\|=2/\xi\gg 1$ and this ratio increases as $a^{1/2}$.
4.2.2 Solitons
Figure 12: Test of an isolated soliton. (Left) Agreement between the initial and final density profiles after $5\,T_{\rm s,phase}$. (Right) Evolution of the wave function at the soliton center. The central density remains approximately constant, while the real and imaginary parts oscillate with a period consistent with Eq. (80)
Solitons are spherically symmetric, non-dispersive, ground-state solutions of the self-gravitating Schrödinger–Poisson system (see Section 1.2.2). Their density profiles feature a flat core supported by quantum pressure, followed by a steep monotonic decrease (see Fig. 12). Furthermore, these solutions obey the scaling symmetry described in Eq. (25).
The soliton solutions can be obtained by inserting the stationary ansatz $\psi(r,t)=e^{-i\omega t}\Psi(r)$ into Eqs. (12), where $\omega$ is the angular frequency and $\Psi$ is real. This form implies a coherent phase and zero bulk velocity. The substitution yields a coupled system of second-order ordinary differential equations, which can be solved numerically with appropriate boundary conditions by seeking the lowest-energy, node-free solution (Guzmán and Ureña-López 2004; Marsh and Pop 2015). Although no exact analytical solution exists, numerical solutions are well approximated by Eq. (23) within $\sim 3\,r_{\rm s}$, enclosing approximately $95\%$ of the total soliton mass (Schive et al. 2014a).
This setup is particularly suited for validating numerical accuracy in the fully nonlinear regime. As an illustration, Fig. 12 shows simulation results obtained using the global Fourier method over a domain of $10\,r_{\rm s}$ with spatial resolution $\Delta h\approx 0.08\,r_{\rm s}$. The simulation is evolved for $5\,T_{\rm s,phase}$, where
$T_{\rm s,phase}\approx 38.2\left(\frac{\rho_{\rm s,max}}{\,\rm{M}_{\odot}\,\rm pc^{-3}}\right)^{-1/2}\,\rm Myr$ (80)
is the oscillation period of the soliton phase, with $\rho_{\rm s,max}$ denoting the maximum soliton density (Guzmán and Ureña-López 2004; Veltmaat et al. 2018; Chiang et al. 2021). The initial and final density profiles match closely. The central density exhibits small-amplitude oscillations, primarily due to the limited accuracy of the fitting function Eq. (23) beyond $3\,r_{\rm s}$. The phase of the wave function oscillates with a period consistent with Eq. (80).
In addition to testing a stationary soliton solution, one can also simulate a perturbed configuration. In such cases, the central density undergoes periodic oscillations with a characteristic timescale given by (Guzmán and Ureña-López 2004; Veltmaat et al. 2018; Chiang et al. 2021)
$T_{\rm s,density}\approx 92.1\left(\frac{\rho_{\rm s,max}}{\,\rm{M}_{\odot}\,\rm pc^{-3}}\right)^{-1/2}\,\rm Myr\approx 2.4\,T_{\rm s,phase}.$ (81)
A perturbed soliton with an oscillation amplitude of order unity is commonly observed in soliton–halo systems forming in cosmological simulations. This behavior can be interpreted as the superposition of the ground-state soliton and excited states (Li et al. 2021).
The stable soliton configuration offers a convenient simulation testbed for exploring various properties of FDM. Example applications include studies of soliton–halo systems through soliton mergers (Schive et al. 2014b; Schwabe et al. 2016; Mocz et al. 2017; Li et al. 2021; Chan et al. 2024; Stallovits and Rindler-Daller 2025; Blum et al. 2025), tidal stripping of solitons (Du et al. 2018), dynamical interactions between solitons and stars (Chan et al. 2018; Chiang et al. 2021), dynamical friction of massive objects within solitons (Wang and Easther 2022; Boey et al. 2024, 2025), modeling the central molecular zone of the Milky Way (Li et al. 2020), and investigating the enhanced growth of supermassive black holes at high redshifts (Chiu et al. 2025).
4.2.3 Isolated halos
An isolated halo serves as a stringent benchmark for testing FDM simulations. A robust FDM algorithm, whether wave- or fluid-based, must meet the following criteria. The halo density profile within the virial radius should be stable on a dynamical timescale. The total mass, momentum, and energy must be conserved with errors below the few-percent level. The kinetic and thermal energies (Eqs. 810) are expected to exhibit energy equipartition, and their sum should follow the virial condition. The density granules need to be well resolved throughout the entire halo.
The central soliton core must also be well resolved and remain stable. In a typical system, the soliton peak density exceeds its ambient halo density by more than an order of magnitude (Schive et al. 2014a, b), with a sharp soliton–halo transition at $r\approx 3.3\text{--}3.5\,r_{\rm s}$ (Mocz et al. 2017; Chiang et al. 2021). Within this radius, the density profile should closely follow Eq. (23) and be supported by quantum pressure, with negligible turbulent motion. The soliton peak density is expected to oscillate with an amplitude of order unity and a characteristic period given by Eq. (81). In addition, the soliton undergoes a confined random walk with respect to the halo center of mass, with a characteristic displacement and timescale on the order of $r_{\rm s}$ and $T_{\rm s,phase}$, respectively (Schive et al. 2020; Li et al. 2021; Dutta Chowdhury et al. 2021). Over longer timescales comparable to the condensation time, the soliton mass may exhibit secular growth (Levkov et al. 2018; Eggemeier and Niemeyer 2019; Chen et al. 2021; Chan et al. 2024).
The isolated halo test provides a valuable means of calibrating the resolution requirements for an FDM algorithm in a fully nonlinear, self-gravitating system. Specifically, it can be used to determine the optimal spatial and temporal resolution needed to resolve the characteristic length and time scales of the ubiquitous density granules, as inferred from Eqs. (2627). In principle, these resolution criteria are broadly applicable to other FDM simulations. An additional complication arises for Lagrangian approaches due to their inherently fixed mass resolution—they may lack sufficient spatial resolution to resolve density granules and quantum pressure in the low-density halo outskirts (see Section 3.1). Furthermore, the phase discontinuity and the divergence of velocity and quantum potential near vortices could present significant numerical challenges for fluid-based schemes (see Section 3.2). It also remains an open question whether fluid schemes can accurately capture the soliton core in a soliton–halo system, characterized by a high-density plateau, a sharp soliton–halo transition, and negligible velocity dispersion (Nori and Baldi 2021; Nori et al. 2023; Chan et al. 2025).
Figure 13: Test of an isolated halo with $M_{\rm h}=1.9\times 10^{10}\,\rm{M}_{\odot}$ and $m_{22}=0.8$. (Left) Density profiles at the beginning and after evolving for approximately one free-fall time, demonstrating halo stability. The central core aligns well with the soliton solution, while the outer region follows the NFW model. (Right) Velocity profiles. Outside the soliton, the thermal velocity $w$ and bulk velocity $v$ coincide due to energy equipartition, and both follow the profile derived from the isotropic spherical Jeans equation (dash-dotted line). Inside the soliton (yellow shaded region), the thermal velocity dominates, and its average value $\left\langle{w}\right\rangle_{s}$ matches that of the inner halo (blue shaded region), indicating thermal equilibrium. See also Fig. 14
As a demonstration, Fig. 13 shows the density and velocity profiles of an FDM halo extracted from a cosmological simulation with $m_{22}=0.8$ at $z=0$. We extract a cubic region of width $186\,\rm kpc$, centered on a halo with a virial radius $r_{\rm h}=72\,\rm kpc$ and a mass $M_{\rm h}=1.9\times 10^{10}\,\rm{M}_{\odot}$ (see Data availability for the link to download the halo data). The central soliton has a radius $r_{\rm s}=0.52\,\rm kpc$ and a mass $M_{\rm s}=1.5\times 10^{8}\,\rm{M}_{\odot}$, with a soliton–halo transition approximately at $4\,r_{\rm s}$. The halo density profile at $r\gtrsim 8\,r_{\rm s}$ is well fitted by the NFW model with a concentration parameter of $c=5.4$. To validate the system’s stability, we evolve it for another $2.3\,\rm Gyr$ (roughly one free-fall time), using the global Fourier method employed in the GAMER code on a $N=1024^{3}$ grid without AMR, corresponding to a spatial resolution of $0.18\,\rm kpc$. We apply isolated boundary conditions for the Poisson solver and periodic boundary conditions for the wave function.
The left panel demonstrates that the density profile outside the soliton remains stable throughout the evolution. The central core closely follows the soliton solution Eq. (23), with minor differences between the initial and final profiles attributable to stochastic soliton oscillations. The right panel shows that the thermal and bulk velocities in the inner halo are closely aligned, indicating energy equipartition. Both velocities match well the velocity dispersion predicted by the isotropic spherical Jeans equation (Dutta Chowdhury et al. 2021). Furthermore, within the soliton, the bulk velocity is negligible, while the average thermal velocity $\left\langle{w}\right\rangle_{\rm s}$ approximately matches the inner-halo thermal velocity, suggesting thermal equilibrium between the soliton and its ambient halo. See Liao et al. (2025) for further discussion.
Figure 14: Test of the same isolated halo shown in Fig. 13. The panels show slices of mass density $\rho$ (upper left), bulk velocity $v$ (upper right), stress tensor $\Sigma_{yy}$ (lower left), and quantum potential $Q$ (lower right) through the halo center. $\rho$ and $\Sigma_{yy}$ peak at the central soliton, whereas $v$ and $Q$ display filamentary structures and reach extreme values at isolated points, corresponding to vortices in density voids. The solid circle indicates the halo virial radius
Fig. 14 shows the mass density $\rho$, bulk velocity magnitude $v$, quantum stress tensor $\Sigma_{yy}$ (see Eq. 18), and quantum potential $Q$ (see Eq. 14) on a slice through the soliton center. Both $\rho$ and $\Sigma_{yy}$ exhibit relatively smooth distributions and peak at the central soliton. By contrast, $v$ and $Q$ show filamentary structures that trace low-density voids caused by destructive interference. Moreover, there are isolated points with extremely high values of $v$ and $Q$, corresponding to locations where vortex lines intersect the plotting plane (see Section 3.2).
In addition to being extracted from cosmological simulations, isolated FDM halos can also be constructed from scratch via the gravitational collapse of a bound system (e.g., Li et al. 2021; Hui et al. 2021; Blum et al. 2025) or from a prescribed distribution function (see Section 2.5). Furthermore, beyond their use in calibrating FDM algorithms, isolated halos serve as versatile simulation testbeds for investigating several distinctive FDM phenomena, such as the formation and dynamics of quantized vortices (Hui et al. 2021), the stochastic motion of nuclear objects (Schive et al. 2020; Dutta Chowdhury et al. 2021; Li et al. 2021), and dynamical heating in dwarf galaxies (Dalal and Kravtsov 2022; Dutta Chowdhury et al. 2023; Teodori et al. 2026), galactic disks (Yang et al. 2024), and stellar streams (Dalal et al. 2021).
4.2.4 Cosmological simulations
Cosmological simulations arguably provide the most comprehensive test for FDM algorithms. For wave-based schemes, sufficient resolution to resolve the de Broglie wavelength must be maintained throughout the entire computational domain. Simulations lacking adequate resolution in large-scale, low-density regions outside halos may misestimate the flow velocity, thereby biasing halo formation and merger histories. Within halos, insufficient resolution may lead to errors in the quantum pressure and turbulence, resulting in incorrect density profiles, soliton properties, and dynamical heating effects from density granules. See Section 3.1 for related discussion. In addition, mass conservation error provides a useful diagnostic of numerical accuracy for non-conservative numerical schemes, such as wave-based finite-difference and local pseudo-spectral methods. For AMR, these resolution requirements help calibrate grid refinement criteria, which are nontrivial and depend on the adopted evolution schemes. Furthermore, numerical artifacts at coarse–fine interfaces must be carefully assessed and minimized.
Fluid-based schemes can accurately capture large-scale structure with reduced resolution requirements, as they do not need to resolve the de Broglie wavelength. However, as discussed in Sections 3.1 and 4.2.3, it remains debated whether these schemes can reliably capture substructures within FDM halos, such as density granules, vortices, and solitons. For hybrid approaches, it is essential to monitor and minimize numerical artifacts at fluid–wave interfaces, ensure consistency with high-resolution wave-only simulations, and demonstrate superior performance compared to those simulations.
Validation of numerical convergence with increasing resolution can be challenging, and sometimes prohibitively expensive, for FDM cosmological simulations, especially for wave schemes with higher $m_{22}$ or larger simulation volumes. This difficulty arises primarily from the need to (i) use smaller time steps ($\Delta t\propto 1/m_{22}\sigma_{\rm 1D}^{2}$) and (ii) apply higher spatial resolution ($\Delta h\propto 1/m_{22}\sigma_{\rm 1D}$) across a larger fraction of the simulation domain than in CDM simulations. An added complication is the potential formation of spurious halos (see Section 3.3). A practical and efficient alternative for assessing numerical convergence is to perform collisionless NN-body simulations using the same initial conditions (Liao et al. 2025). First, the halo masses and the density profiles outside the central solitons in the genuine FDM and NN-body simulations should agree closely. Note that a good fit to an NFW profile beyond the soliton does not guarantee numerical accuracy, as an incorrect profile may still match an NFW model but with erroneous halo mass and concentration. Second, the halo positions should be consistent across the two types of simulations, since the spatial distribution of spurious halos is primarily driven by numerical noise (see Fig. 8).
To obtain the initial conditions for FDM cosmological simulations, one can first use the Boltzmann code axionCAMB to compute the FDM initial power spectrum. Subsequently, one can use an initial condition tool such as MUSIC (Hahn and Abel 2011), N-GenIC (Springel 2015), or MPgrafic (Prunet et al. 2008) to construct the three-dimensional comoving mass density $\tilde{\rho}$ and comoving peculiar velocity $\bm{\tilde{v}}=a\bm{v_{\rm pec}}$. For wave-based schemes and fluid-based schemes that evolve the phase field, the comoving wave function is given by $\tilde{\psi}=\tilde{\rho}^{1/2}e^{i\tilde{S}}$, where $\bm{\tilde{v}}=(\hbar/m)\bm{\tilde{\nabla}}\tilde{S}$ (see Section 1.1). Accordingly, the initial wave function can be constructed by solving
$\|\tilde{\psi}\|$ $=\tilde{\rho}^{1/2},$ (82)
$\tilde{\nabla}^{2}\tilde{S}$ $=\frac{m}{\hbar}\bm{\tilde{\nabla}}\cdot\bm{\tilde{v}}$. (83)
Figure 15: Cosmological simulation in a comoving box of size $L=4\,h^{-1}\rm Mpc$ with $m_{22}=0.1$ at $z=0$. (Left) Projected density map centered on a halo of mass $M_{\rm h}=5.9\times 10^{11}\,h^{-1}\rm{M}_{\odot}$. (Right) AMR grid map showing the regions that employ the fluid and wave schemes, respectively, in the hybrid algorithm of Kunkel et al. (2025)
As an illustration, Figs. 15 and 16 show results from FDM cosmological simulations with $m_{22}=0.1$ at $z=0$ in a comoving box of size $L=4\,h^{-1}\rm Mpc$. See Data availability for the link to download the initial condition files. Fig. 15 shows the projected mass density and AMR grid distribution from a simulation using the GAMER code with a hybrid scheme (Kunkel et al. 2025). This simulation employs a $64^{3}$ root grid and 8 additional refinement levels, achieving a maximum resolution of $0.24\,h^{-1}\rm kpc$. The fluid scheme that solves the Hamilton–Jacobi–Madelung equations is applied at levels $0\text{--}3$, which cover smooth, low-density regions and occupy the majority of the simulation volume. In comparison, the wave scheme based on the FC–Gram algorithm is applied at levels $4\text{--}8$, which target regions with strong interference, such as filaments and halos, and dominate the computational cost.
Figure 16: Density profiles of the central halo in Fig. 15 at $z=0$ obtained with different evolution schemes from the same initial condition. For the FDM simulation using a hybrid scheme with sufficient resolution (solid line), the central core is well fitted by the soliton solution (thin dotted lines), and the outer profile matches the collisionless NN-body simulation (thick dotted line). When the same hybrid scheme is applied but with adequate wave-scheme resolution only inside the soliton (dashed line), both the central soliton and the outer profile become overly concentrated, although the halo mass remains accurate, as indicated by the corresponding halo virial radius (vertical lines). In contrast, using the global Fourier method with insufficient resolution to resolve the de Broglie wavelength severely underestimates the halo mass (dash-dotted line)
Fig. 16 compares the density profiles of the central halo in Fig. 15 at $z=0$ obtained with different evolution schemes. This halo has a mass $M_{\rm h}=5.9\times 10^{11}\,h^{-1}\rm{M}_{\odot}$ and a virial radius $r_{\rm h}=166\,h^{-1}\rm kpc$. The central soliton has a mass $M_{\rm s}=2.2\times 10^{9}\,h^{-1}\rm{M}_{\odot}$ and a radius $r_{\rm s}=1.1\,h^{-1}\rm kpc$. The free-fall velocity outside the halo is approximately $110\,\rm km\,\rm s^{-1}$, corresponding to a de Broglie wavelength of $\lambda_{\rm dB,ff}\sim 7.6\,h^{-1}\rm kpc$. For the hybrid scheme with fiducial resolution, the central profile aligns well with the soliton solution (Eq. 23), and the outer profile closely matches that of a collisionless NN-body simulation using the GADGET-2 code with the same initial condition. In contrast, the global Fourier method, performed with a $1024^{3}$ grid and $3.9\,h^{-1}\rm kpc$ spatial resolution, underestimates the halo mass because the resolution is insufficient to properly resolve $\lambda_{\rm dB,ff}$. When the same hybrid scheme is applied with adequate resolution confined to the soliton but inadequate resolution elsewhere in the halo, quantum pressure and turbulence are underestimated, leading to unphysical halo contraction. The deepened gravitational potential raises the halo temperature and, in turn, increases the soliton energy and mass. Nevertheless, note that (i) the resulting central profile still fits a soliton solution, and (ii) the halo mass remains in good agreement with the fiducial value. The latter is attributed to the fluid scheme employed outside the halo. These observations suggest that agreement in halo mass and a good match to the theoretical soliton profile are necessary but not sufficient to establish numerical accuracy (Liao et al. 2025).
Due to the extremely high computational cost, most FDM cosmological simulations to date are limited to probing relatively small FDM particle masses ($m_{22}\sim 0.1\text{--}2$)—either in small simulation boxes ($L\sim 1\text{--}2\,h^{-1}\rm Mpc$) evolved to $z\sim 0\text{--}1$, or in somewhat larger boxes ($L\sim 5\text{--}20\,h^{-1}\rm Mpc$) but only to higher redshifts ($z\gtrsim 3$). Representative examples include simulations using the global Fourier method, applied to either pure FDM scenarios (Woo and Chiueh 2009; Li et al. 2019; May and Springel 2021, 2022; Chan et al. 2022) or a mixed CDM–FDM model (Laguë et al. 2024), and the wave-based finite-difference scheme with AMR (Schive et al. 2014a, b). Simulations employing hybrid schemes and zoom-in techniques enable resolving more massive halos, simulating larger volumes, or reaching lower redshifts (Veltmaat et al. 2018; Schwabe and Niemeyer 2022; Kunkel et al. 2025; Chan et al. 2025; Chiu et al. 2025). In comparison, NN-body (Schive et al. 2016; Sarkar et al. 2016; Iršič et al. 2017; Corasaniti et al. 2017; Armengaud et al. 2017; Kobayashi et al. 2017; Ni et al. 2019; Leong et al. 2019; Nadler et al. 2025) and SPH (Zhang et al. 2018a; Nori et al. 2019; Nori and Baldi 2021; Nori et al. 2023) simulations can explore much higher $m_{22}$ and larger volumes at lower redshifts, albeit each with its own limitations. A few pioneering simulations have incorporated baryonic physics (Mocz et al. 2019, 2020; Veltmaat et al. 2020), though they are currently restricted to $m_{22}=2.5$, $L\sim 2\,h^{-1}\rm Mpc$, and $z\sim 4\text{--}5$.
5 Conclusions
In this review, we introduced a variety of FDM algorithms (Section 2), described the associated numerical challenges (Section 3), and presented a representative set of numerical tests (Section 4). Wave-based methods most accurately capture the fine-grained wave phenomena in FDM halos, such as density granulation, vortices, and soliton cores (Section 1.2). However, they are computationally expensive due to the need to resolve the de Broglie wavelength and its rapid oscillations. Moreover, they do not, in general, guarantee conservation of momentum and energy.
Fluid-based methods are significantly more efficient for modeling smooth, high-velocity flows. They ensure manifest conservation and can readily adopt a Lagrangian formulation that preserves Galilean invariance. However, they struggle to resolve wave features in interference-dominated, multi-stream regions. Hybrid schemes, coupled with AMR, apply fluid methods in smooth, single-stream regions and switch to wave methods in complex, multi-stream regions, thereby enabling simulations of larger volumes and higher FDM particle mass. Nevertheless, simulating halos more massive than $\sim 10^{12}\,\rm{M}_{\odot}$ to $z\sim 0$ with $m_{22}\gg 1$ likely remains computationally infeasible in the near future.
Several alternative approaches are also worth noting. Eigenmode methods offer efficient tools for constructing FDM halos or simulating restricted halo regions, provided that gravitational backreaction from non-dark-matter components can be neglected. Collisionless NN-body methods, while neglecting quantum pressure, remain valuable for probing large-scale structure and assessing the numerical convergence of genuine FDM cosmological simulations. Finally, machine learning, though not addressed in this review, offers an interesting direction for accelerating or augmenting FDM simulations (e.g., Mishra and Tolley 2025).
The discussions in this review have mostly focused on a single FDM species without self-interaction. However, the algorithms described here can be easily extended to explore more general scenarios—for example, FDM with self-interaction (e.g., Mocz et al. 2023; Jain and Amin 2023; Painter et al. 2024; Glennon et al. 2024; Stallovits and Rindler-Daller 2025), vector or multi-component FDM (e.g., Amin et al. 2022; Huang et al. 2023; Gosenca et al. 2023; Luu et al. 2024), and mixed CDM–FDM models (e.g., Schwabe et al. 2020; Laguë et al. 2024).
To facilitate comparison among different FDM codes, we provide links to the initial condition files for the isolated-halo and cosmological simulations in Data availability.
Data availability
The initial condition files of the numerical tests presented in Sections 4.2.3 (isolated halos) and 4.2.4 (cosmological simulations) can be downloaded from the link: https://drive.google.com/drive/folders/1Xjgi6AYtl8aNhqLplPhnt_FVSooyBQeA?usp=sharing.
[…]
Acknowledgments
I would like to thank Pin-Yu Liao for polishing the figures and Hsinhao Huang, Alexander Kunkel, and Guan-Ming Su for their helpful comments. We use yt (Turk et al. 2011) for data visualization and analysis. This research is partially supported by the National Science and Technology Council (NSTC) of Taiwan under Grant No. NSTC 111-2628-M-002-005-MY4 and the NTU Academic Research–Career Development Project under Grant No. NTU-CDP-113L7729.
Declarations
[…]
Conflict of interest The author declares no conflict of interest.